### v1

# Polycyclic codes over serial rings and their annihilator CSS construction

Abstract
In this paper, we investigate the algebraic structure for polycyclic codes over a specific class of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ for $i\in\{1,\ldots,s\}$ is a monic square-free polynomial. We define quasi-$s$-dimensional polycyclic codes and establish an $R$-isomorphism between these codes and polycyclic codes over $\mathscr{R}$. We provide necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator linear complementary dual, and annihilator dual-containing polycyclic codes over this class of rings. We also establish the CSS construction for annihilator dual-preserving polycyclic codes over the chain ring $R$ and use this construction to derive quantum codes from polycyclic codes over $\mathscr{R}$.
keywords:
Polycyclic code, Serial ring, Annihilator dual, quasi-$s$-dimensional polycyclic code, annihilator CSS construction AMS Subject Classification 2010: 94B15, 13B25, 81P70
journal: Cryptography and Communicationsmmfootnotetext: Maryam Bajalan is supported by the Bulgarian Ministry of Education and Science, Scientific Programme ”Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No. DO1-67/05.05.2022.vvfootnotetext: Both authors were partially supported by Grant TED2021-130358B-I00 funded by MCIU/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”
1 Introduction
Polycyclic codes are ideals in the quotient of the polynomial ring $A[x]/\langle f(x)\rangle$, where $A$ is a ring and $f(x)$ is a polynomial in $A[x]$, see [1, 5, 6, 23]. Cyclic codes and constacyclic codes over $A$ are particular instances when $f(x)=x^{n}-1$ and $f(x)=x-\lambda$, respectively. A ring is called uniserial if all its submodules are linearly ordered by inclusion, and it is called serial if it is a direct sum of uniserial rings. In this paper, we aim to investigate polycyclic codes in the case $A$ is the serial ring $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial for $i=1,2,\ldots,s$.
There is a vast literature of polycyclic codes over some special classes of rings, that we slightly review here. Alkenani et al. [3] studyed constacyclic codes over the ring $\mathbb{F}_{q}[u_{1},u_{2}]/\langle u_{1}^{2}-u_{1},u_{2}^{2}-u_{2},u_{1}u_{2}-u_{2}u_{1}\rangle$. Dinh et al. [10] investigated constacyclic codes over the ring $\mathbb{F}_{p}+u_{1}\mathbb{F}_{p}+\cdots+u_{s}\mathbb{F}_{p}$, subject to the conditions $u_{i}^{2}=u_{i}$, $u_{i}u_{j}=u_{j}u_{i}=0$, and $p$ is an odd prime. Dinh et al. [12] explored the algebraic structure of constacyclic codes over finite semi-simple rings, defined as direct sums of finite fields. Islam et al. [21] focused their study on cyclic codes over the non-chain ring $\mathbb{F}_{q}[u]/\langle u^{e}-1\rangle$. Bag et.al [4] examined skew constacyclic codes over the ring $\mathbb{F}_{q}[u,v]/\langle u^{2}-1,v^{2}-1,uv-vu\rangle$. Ali et al. [2] worked out the case cyclic codes over non-chain rings $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{s}]/\langle u^{2}_{i}-(\alpha_{i})^{2},\quad u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $\alpha_{i}$ is a non-zero element of $\mathbb{F}_{q}$. Goyal and Raka [15, 16] explored polyadic cyclic and constacyclic codes over the non-chain ring $\mathbb{F}_{q}[u,v]/\langle f(u),g(v),uv-vu\rangle$, with $f(u)$ and $g(v)$ splitting into distinct linear factors over $\mathbb{F}_{q}$. Bhardwaj et al. [7] investigated constacyclic codes over the finite non-chain ring $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{k}]/\langle f_{i}(u_{i}),u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $f_{i}(u_{i})$ is a polynomials that splits into distinct linear factors over $\mathbb{F}_{q}$. Qi [32] focused on polycyclic codes over $\mathbb{F}_{q}+u\mathbb{F}_{q}$, with $u^{2}=0$. Karthick [22] enquired about polycyclic codes over the ring $R=\mathbb{F}_{q}+u\mathbb{F}_{q}+v\mathbb{F}_{q}$, subject to the conditions $u^{2}=\alpha u$, $v^{2}=v$, and $uv=vu=0$, where $\alpha$ represents a unit element of $R$.
The serial ring $\mathscr{R}=\frac{R[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}$, where $R$ is a finite commutative chain ring with maximal ideal $\mathfrak{m}$ and the natural projection of $t_{i}$ on $(R/\mathfrak{m})[x_{i}]$ is a square free polynomial for each $i\in\{1,\ldots,s\}$, is a comprehensive generalization of the rings previously discussed in the papers [2, 3, 4, 7, 10, 15, 16, 21, 22, 32]. Note that $R$ is a Frobenius ring and when $R$ is assumed to be a field the serial ring $\mathscr{R}$ is a semi-simple ring, see Proposition 2.
In this paper, we investigate polycyclic codes over the general ring $\mathscr{R}$. The outline of the paper is as follows. Section 2 shows the foundational structure of $\mathscr{R}$. In Section 3, we examine how a polycyclic code $\mathscr{C}$ over $\mathscr{R}$ decomposes into a direct sum $\oplus_{i=1}^{k}\mathscr{C}_{i}$, where each $\mathscr{C}_{i}$ constitutes a polycyclic code over a chain ring. Section 4 explores the concept of quasi-2-dimensional polycyclic (Q2DP) codes along with their multidimensional counterparts (QsDP), which extend the ideas of quasi-polycyclic codes and multidimensional cyclic codes [18, 19, 20, 27, 36]. Furthermore, we define the concept of $f(x)$-polycyclic-QsDP codes as a specific subclass of QsDP codes. Then, we prove that when considering polycyclic codes over $\mathscr{R}$ as $R$-submodules, they precisely coincide with $f(x)$-polycyclic-QsDP codes. Polycyclic codes are defined as the invariant $\mathscr{R}$-submodules of $\mathscr{R}^{n}$ under the operator $\tau_{f(x)}$. Their counterpart are sequential codes, which are defined as invariant subspace under the adjoint operator of $\tau_{f(x)}$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator. Consequently, a polycyclic code $\mathscr{C}$ and its Euclidean dual $\mathscr{C}^{\perp}$ do not share the same structure. Thus, the first ones are equivalent to ideals in $\mathscr{R}[x]/\langle f(x)\rangle$, defining a polynomial and ideal structure for the second ones is challenging. Note that, if $\mathscr{C}$ is an ideal in $\mathscr{R}[x]/\langle f(x)\rangle$ generated by the polynomial $g(x)$ with the property that $f(x)=g(x)h(x)$, it cannot be assumed that $\mathscr{C}^{\perp}$ is generated by $h(x)$ because $\mathscr{C}^{\perp}$ is not necessarily a polycyclic code. This incorrect interpretation happened in the proof of Theorem 5 in [30]. In Section 5, using multivariable codes over chain rings in [24, 25], we establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case where $f(x)$ is a polynomial with coefficients in $R$. The annihilator dual is an alternative duality concept ensuring that the dual of polycyclic codes has a polycyclic structure, a feature that fails in Euclidean dual. In Section 6, some properties of the annihilator dual of a polycyclic code over $\mathscr{R}$ are studied. Usually, it is stated in the literature that “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=g^{2}(x)$. As explained in Remark 5, that statement relies on the assumption that “if $b$ divides $c$ and $c$ divides $b$, then $b$ equals $c$’, without considering that the divisibility relation for rings is different, namely “if $b$ divides $c$ and $c$ divides $b$, then $\langle b\rangle$ equals $\langle c\rangle$. Thus in Section 6, we revise this property as follows, “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=ag^{2}(x)$ for some unit element $a$ in the ring $\mathscr{R}$. To finish this paper, we study some quantum code constructions from this type of codes. Calderbank-Shor-Steane (CSS) construction allows for constructing quantum codes from dual-preserving classical codes, see for example [29, 33]. Existing literature has applied this construction only to codes that preserve Euclidean or Hermitian duality. In this paper, we present construction based on preserving the annihilator duality. In Section 7, we provide an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.
2 The structure of the ring $\mathscr{R}=\frac{R[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}$
Let $R$ be a finite commutative chain ring, and let $\gamma\in R$ be a fixed generator of the maximal ideal $\mathfrak{m}$ with nilpotency index $e$. Consider $\mathbb{F}_{q}=R/\mathfrak{m}$ as the residue field, where $q=p^{t}$, $p$ a prime number. Let ¯$\overline{\phantom{x}}:R[y]\to\mathbb{F}_{q}[y]$ represent the natural ring homomorphism, mapping $r\in R$ to $r+\mathfrak{m}$ and the variable $y$ to itself. A polynomial $p(y)=\sum_{i=0}^{n-1}p_{i}y^{i}\in R[y]$ is called square-free (basic irreducible) if $\bar{p}(y)$ is square-free (basic irreducible) in $\mathbb{F}_{q}[y]$. Additionally, it is called regular if it is not a zero divisor in $R[y]$. In [11, Proposition 2.3] it is stated that $p(y)$ is regular if and only if $p_{i}$ is a unit element for some $i$. Hence, monic polynomials are regular. If $g(y)\in R[y]$ is a regular basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ forms a chain ring, see [11, Lemma 3.1]. It is well-known that the monic square-free polynomial $g(y)$ factors uniquely as a product of monic basic irreducible pairwise coprime polynomials, given by $g(y)=g_{1}(y)g_{2}(y)\cdots g_{r}(y)$[28, Theorem XIII.4]. Hence, by the Chinese Remainder Theorem, the quotient ring $R[y]/\langle g(y)\rangle$ can be decomposed as a direct sum of chain rings
$\frac{R[y]}{\langle g(y)\rangle}\cong\frac{R[y]}{\langle g_{1}(y)\rangle}\bigoplus\frac{R[y]}{\langle g_{2}(y)\rangle}\bigoplus\cdots\bigoplus\frac{R[y]}{\langle g_{r}(y)\rangle}\cdot$ (1)
The algebraic properties of the algebra $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, assuming that each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial of degree $m_{i}$, was explored in [24]. Let $H_{i}$ denote the set of roots of $\bar{t}_{i}(x_{i})$ in a suitable extension of $\mathbb{F}_{q}$. For each $\nu\in\mathcal{H}=\prod_{i=1}^{s}H_{i}$, let $C(\nu)$ denote the set of the class of $\nu$ as $\{(\nu^{q^{j}_{1}},\ldots,\nu^{q^{j}_{s}})\mid j\in\mathbb{N}\}$. The set of all these classes, denoted by $\mathcal{C}=\mathcal{C}(t_{1},\ldots,t_{s})$, forms a partition of $\mathcal{H}$. The size of each class is given by $|C(\nu)|=\text{lcm}(d_{1},\ldots,d_{s})=[\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s}):\mathbb{F}_{q}]$, where $d_{i}$ is the degree of the irreducible polynomial of $\nu_{i}$ over $\mathbb{F}_{q}$. Let $p_{C,i}(x_{i})$ denote the polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$ and $C=C(\nu)$ is a class of $\nu$. For each $i=2,\ldots,s$, let $b_{C,i}(x_{i})=\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))\in\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1})[x_{i}]$ and $\widetilde{b}_{C,i}(x_{i})=\frac{p_{C,i}(x_{i})}{b_{C,i}(x_{i})}$. The polynomials $b_{C,i}(x_{i})$ and $\widetilde{b}_{C,i}(x_{i})$ are coprime polynomials. One can define the multivariable polynomials $w_{C,i}(x_{1},\ldots,x_{i})$ and $\pi_{C,i}(x_{1},\ldots,x_{i})$ obtained by substituting $\nu_{i}$ by $x_{i}$ in polynomials $b_{C,i}(X_{i})$ and $\widetilde{b}_{C,i}(x_{i})$, respectively and denote the Hensel lifts to $R$ of the polynomials $p_{C,i}$, $w_{C,i}$, and $\pi_{C,i}$ by $q_{C,i}$, $z_{C,i}$, and $\sigma_{C,i}$, respectively. If we denote $\langle q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle$ by $I_{C}$, then the ring $T_{C}=R[x_{1},\ldots,x_{s}]/I_{C}$ is a chain ring with maximal ideal $\mathfrak{M}=\langle a,q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle+I_{C}$, and the residue field $T_{C}/\mathfrak{M}\cong\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s})$. Let us define the polynomial
$h_{C}(x_{1},x_{2},\ldots,x_{s}):=\prod_{i=1}^{s}t_{i}(x_{i})q_{C,i}(x_{i})\prod_{i=1}^{s}\sigma_{q_{C,i}}(x_{2},\ldots,x_{i}).$ (2)
The following two propositions outline the properties of the serial ring $\mathscr{R}$.
Proposition 1 ([24]).
  1. 1.
    $\mathscr{R}\cong\bigoplus_{C\in\mathcal{C}}\langle h_{C}+I\rangle\cong\bigoplus_{C\in\mathcal{C}}R[x_{1},\ldots,x_{s}]/I_{C}$, where $\mathcal{C}$ is the partition of $\mathcal{H}$.
  2. 2.
    $R[x_{1},\ldots,x_{s}]/I_{C}$ is a chain ring with maximal ideal $\mathfrak{M}=\langle a,q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle+I_{C}$ and its residue field is given by $T_{C}/\mathfrak{M}\cong\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s})$.
  3. 3.
    There exists a complete set of orthogonal idempotent elements $\{e_{C}\in\mathscr{R}\mid C\in\mathcal{C}\}$ such that $\mathscr{R}e_{C}\cong\langle h_{C}+I\rangle$.
  4. 4.
    If $\lvert\mathcal{C}\rvert=k$ and $\{C_{i}\}_{i=1}^{k}$ is a set of representatives $\mathcal{C}$, then the ring $\mathscr{R}$ can be decomposed as a direct sum
    $\mathscr{R}\cong\bigoplus_{i=1}^{k}\mathscr{R}e_{C_{i}}\cong\bigoplus_{i=1}^{k}\mathscr{R}_{i},$
    where each $\mathscr{R}_{i}$ is a chain ring. Moreover, any element $r\in\mathscr{R}$ can be expressed uniquely as $r=r_{1}e_{C_{1}}+r_{2}e_{C_{2}}+\cdots+r_{k}e_{C_{k}}$, where $r_{i}\in\mathscr{R}$.
A ring $A$ is called uniserial if all $A$-submodules of the module $A_{A}$ are linearly ordered by inclusion. A ring is called serial if it is a direct sum of uniserial rings. Any uniserial ring is local, and a serial ring is uniserial if and only if it is local, see [31, Section 1.4]. Note that any chain ring is a serial ring, thus, if $g(y)\in R[y]$ is a monic basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ is a serial ring. Finite commutative semi-simple rings are those rings that can be represented as finite direct sums of finite fields. The following proposition is straightforward from the previous paragraph.
Proposition 2.
  1. 1.
    $\mathscr{R}$ is a serial ring.
  2. 2.
    $\mathscr{R}$ is a principal ideal Frobenius ring.
  3. 3.
    If R is a field, then $\mathscr{R}$ is a semi-simple ring.
3 Polycyclic codes over $\mathscr{R}$
A subset $\mathscr{C}\subseteq\mathscr{R}^{n}$ is called a linear code over $\mathscr{R}$ (or an $\mathscr{R}$-linear code) of length $n$ if it is an $\mathscr{R}$-submodule of $\mathscr{R}^{n}$. It is called free code if it is a free $\mathscr{R}$-submodule. Throughout this paper
$f(x)=x^{n}-(f_{n-1}x^{n-1}+\cdots+f_{1}x+f_{0})$ (3)
will denote a monic polynomial in $\mathscr{R}[x]$. An $f(x)$-polycyclic shift $\tau_{f(x)}:\mathscr{R}^{n}\rightarrow\mathscr{R}^{n}$ is defined by
$\tau_{f(x)}(b_{0},b_{1},\ldots,b_{n-1})=(0,b_{0},b_{1},\ldots,b_{n-2})+b_{n-1}(f_{0},f_{1},\ldots,f_{n-1}).$ (4)
An $\mathscr{R}$-linear code of length $n$ is called $f(x)$-polycyclic code if it remains invariant under the $f(x)$-polycyclic shift $\tau_{f(x)}$.
According to Proposition 1, every component $c_{j}\in\mathscr{R}$ in $\boldsymbol{c}=(c_{0},c_{1},\ldots,c_{n-1})\in\mathscr{C}\subseteq\mathscr{R}^{n}$ has a unique representation $c_{j}=c_{1,j}e_{C_{1}}+c_{2,j}e_{C_{2}}+\cdots+c_{k,j}e_{C_{k}}$, where $c_{i,j}\in\mathscr{R}$ for all $i,j$, $1\leqslant i\leqslant k$,  $10\leqslant j\leqslant n-$ 1. Therefore, $\boldsymbol{c}$ can be expressed as
$\boldsymbol{c}=(c_{1,0},c_{1,1},c_{1,2},\ldots,c_{1,n-1})e_{C_{1}}+(c_{2,0},c_{2,1},c_{2,2},\ldots,c_{2,n-1})e_{C_{2}}+\cdots+(c_{k,0},c_{k,1},c_{k,2},\ldots,c_{k,n-1})e_{C_{k}},$ (5)
which we will denote as $\boldsymbol{c}=\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}$. For a code $\mathscr{C}$ of length $n$ over $\mathscr{R}$, we define the subsets $\mathscr{C}_{i}$ of $\mathscr{R}^{n}$ for $1\leqslant i\leqslant k$ as follows:
$\mathscr{C}_{i}=\{\boldsymbol{c}_{i}\in\mathscr{R}^{n}\mid\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}\in\mathscr{C}\}.$
The following result is straightforward.
Theorem 1.
Let $\mathscr{C}$ be a linear code of length $n$ over $\mathscr{R}$. Then, each $\mathscr{C}_{i}$ is a linear code over $\mathscr{R}$ and $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}$.
Similarly to Equation (5), one can rewrite the polycyclic relationship as follows: consider the decomposition by the orthogonal idempotents $\boldsymbol{f}_{i}=(f_{i,0},f_{i,1},\ldots,f_{i,n-1})\in\mathscr{R}^{n}$ with $\boldsymbol{f}=(f_{0},f_{1},\ldots,f_{n-1})\in\mathscr{R}^{n}$, and one can write $\boldsymbol{f}=\boldsymbol{f}_{1}e_{C_{1}}+\boldsymbol{f}_{2}e_{C_{2}}+\cdots+\boldsymbol{f}_{k}e_{C_{k}}.$ Corresponding to each $\boldsymbol{f}_{i}$, we define the polynomial
$f_{i}(x)=x^{n}-(f_{i,n-1}x^{n-1}+\cdots+f_{i,1}x+f_{i,0}).$ (6)
Lemma 1.
Let $\mathscr{C}$ be a linear code over $\mathscr{R}$. Then, $\tau_{f(x)}(\boldsymbol{c})\in\mathscr{C}$ if and only if $\tau_{f_{i}(x)}(\boldsymbol{c}_{i})\in\mathscr{C}_{i}$ for all $1\leqslant i\leqslant k$.
Proof.
Let $\boldsymbol{c}=(c_{0},c_{1},\ldots,c_{n-1})\in\mathscr{C}$. Since $\{e_{C_{1},},\ldots,e_{C_{k}}\}$ is a complete set of orthogonal idempotents, we have
$\tau_{f(x)}(\boldsymbol{c})$ $=(0,c_{0},c_{1},\ldots,c_{n-2})+c_{n-1}(f_{0},f_{1},\ldots,f_{n-1})$
$=\bigg{(}0,\sum_{i=1}^{k}c_{i,0}e_{C_{i}},\sum_{i=1}^{k}c_{i,1}e_{C_{i}},\ldots,\sum_{i=1}^{k}c_{i,n-2}e_{C_{i}}\bigg{)}+\sum_{i=1}^{k}c_{i,n-1}e_{C_{i}}\bigg{(}\sum_{i=1}^{k}f_{i,0}e_{C_{i}},\sum_{i=1}^{k}f_{i,1}e_{C_{i}},\ldots,\sum_{i=1}^{k}f_{i,n-1}e_{C_{i}}\bigg{)}$
$=\bigg{(}0,\sum_{i=1}^{k}c_{i,0}e_{C_{i}},\sum_{i=1}^{k}c_{i,1}e_{C_{i}},\ldots,\sum_{i=1}^{k}c_{i,n-2}e_{C_{i}}\bigg{)}+\bigg{(}\sum_{i=1}^{k}c_{i,n-1}f_{i,0}e_{C_{i}},\sum_{i=1}^{k}c_{i,n-1}f_{i,1}e_{C_{i}},\ldots,\sum_{i=1}^{k}c_{i,n-1}f_{i,n-1}e_{C_{i}}\bigg{)}$
$=\sum_{i=1}^{k}\bigg{(}\big{(}0,c_{i,0},c_{i,1},\ldots,c_{i,n-2}\big{)}+c_{i,n-1}\big{(}f_{i,0},f_{i,1},\ldots,f_{i,n-1}\big{)}\bigg{)}e_{C_{i}}$
$=\sum_{i=1}^{k}\bigg{(}\tau_{f_{i}(x)}(c_{i,0},c_{i,1},\ldots,c_{i,n-1})\bigg{)}e_{C_{i}}=\sum_{i=1}^{k}\bigg{(}\tau_{f_{i}(x)}(\boldsymbol{c}_{i})\bigg{)}e_{C_{i}}.$
Therefore, $\tau_{f(x)}(\boldsymbol{c})\in\mathscr{C}$ if and only if $\tau_{f_{i}(x)}(\boldsymbol{c}_{i})\in\mathscr{C}_{i}$ for all $1\leqslant i\leqslant k$. ∎
One can consider the $\mathscr{R}$-module isomorphism $\phi:\mathscr{R}^{n}\to\mathscr{R}_{f}=\mathscr{R}[x]/\langle f(x)\rangle$, which maps each vector to its corresponding polynomial representation. It was proved in [13] that the image of any $f(x)$-polycyclic code under $\phi$ is an ideal in $\mathscr{R}_{f}$. From now on, we abuse the notation $\mathscr{C}$ and we will also call the ideals in the ring $\mathscr{R}_{f}$ $f(x)$-polycyclic codes. The ring $\mathscr{R}_{f}$ might not always be a principal ideal ring and d a single codeword generator is a nice property for encoding and decoding. Therefore, we determine the condition under which the ring $\mathscr{R}_{f}$ is a principal ideal ring.
Proposition 3.
  1. 1.
    The ring $\mathscr{R}[x]/\langle f(x)\rangle$ is isomorphic to $\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$, where $\mathscr{R}_{i}$ represents the chain ring defined in Proposition 1 and $f_{i}(x)$ denotes the polynomial defined in Equation (6).
  2. 2.
    The ring $\mathscr{R}[x]/\langle f(x)\rangle$ is principal ideal if and only if $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a principal ideal ring for each $i\in\{1,\ldots,k\}$.
Proof.
  1. 1.
    Let $g(x)=g_{0}+g_{1}x+\cdots+g_{n-1}x^{n-1}$ be an element in $\mathscr{R}[x]/\langle f(x)\rangle$. We know that the coefficients $g_{j}\in\mathscr{R}$ are uniquely expressed as $g_{j}=g_{1,j}e_{C_{1}}+g_{2,j}e_{C_{2}}+\cdots+g_{k,j}e_{C_{k}}$, i.e., $g_{j}$ can be uniquely expressed as $(g_{1,j},\,g_{2,j},\,\ldots,\,g_{k,j})$. Therefore, $g(x)=g_{1}(x)e_{C_{1}}+g_{2}(x)e_{C_{2}}+\cdots+g_{k}(x)e_{C_{k}}$, where $g_{i}(x)=g_{i,0}+g_{i,1}x+g_{i,2}x^{2}+\cdots+g_{i,n-1}x^{n-1}\in\mathscr{R}_{i}[x]$. Hence, the polynomial $g(x)$ can also be represented as $k$-tuples of polynomials $g(x)=\big{(}g_{1}(x),g_{2}(x),\ldots,g_{k}(x)\big{)}$. Now, we define the ring isomorphism
    $\eta:\mathscr{R}_{f}=\frac{\mathscr{R}[x]}{\langle f(x)\rangle}$ $\longrightarrow\frac{\mathscr{R}_{1}[x]}{\langle f_{1}(x)\rangle}\times\frac{\mathscr{R}_{2}[x]}{\langle f_{2}(x)\rangle}\times\dots\times\frac{\mathscr{R}_{k}[x]}{\langle f_{k}(x)\rangle},$ (7)
    $g(x)$ $\longmapsto\big{(}g_{1}(x),\,g_{2}(x),\,\dots,\,g_{k}(x)\big{)}.$
  2. 2.
    It is straightforward.
Remark 1.
  • 1.
    The condition for the ring $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ being a principal ideal ring in terms of the chain ring $\mathscr{R}_{i}$ and the polynomial $f_{i}(x)$ can be found in [26, Theorem 1].
  • 2.
    If $f_{i}(x)$ is a regular basic irreducible polynomial in the ring $\mathscr{R}_{i}[x]$, then $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma_{i}+\langle f_{i}(x)\rangle\rangle$, where $\gamma_{i}$ denotes the maximal ideal of the chain ring $\mathscr{R}_{i}$, see [11, Lemma 3.1].
Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$, we will denote the rank of the code $\mathscr{C}$ as $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$, i.e. the minimum number of generators needed to span $\mathscr{C}$ as $\mathscr{R}$-linear combinations. We will also denote the minimum Hamming distance of $\mathscr{C}$ as $d(\mathscr{C})$.
Theorem 2.
Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$. Then we have the following statements.
  1. 1.
    There are $f_{i}(x)$-polycyclic codes $\mathscr{C}_{i}$ in $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ such that $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}$.
  2. 2.
    $\mathscr{C}$ is an $f(x)$-polycyclic code if and only if every $\mathscr{C}_{i}$ is an $f_{i}(x)$-polycyclic code.
  3. 3.
    $\mathscr{C}$ is generated by a single codeword if and only if every $\mathscr{C}_{i}$ is generated by a single codeword.
  4. 4.
    If each $\mathscr{C}_{i}$ is generated by the polynomial $g_{i}(x)$, then $\mathscr{C}$ is generated by $g(x)=\oplus_{i=1}^{k}g_{i}(x)e_{C_{i}}$.
  5. 5.
    $|\mathscr{C}|=\prod_{i=1}^{k}|\mathscr{C}_{i}|$.
  6. 6.
    $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})=\max\{\operatorname{rank}_{\mathscr{R}_{i}}(\mathscr{C}_{i})\mid i=1,\ldots,k\}$.
  7. 7.
    $d(\mathscr{C})=\min\{d(\mathscr{C}_{i})\mid i=1,\ldots,k\}$.
  8. 8.
    $\mathscr{C}$ is a free over $\mathscr{R}$ if and only if every $\mathscr{C}_{i}$ is free over $\mathscr{R}_{i}$ for $i=\{1,\ldots,k\}$.
Proof.
  1. 1.
    The ideal $\mathscr{C}$ in $\mathscr{R}[x]/\langle f(x)\rangle$ can be expressed as $\mathscr{C}=\eta(\mathscr{C}_{1},\mathscr{C}_{2},\ldots,\mathscr{C}_{k})$, where each $\mathscr{C}_{i}$ is an ideal in $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ and $\eta$ is the isomorphism defined in (7). Conversely, the set of ideals $\mathscr{C}_{i}$ in $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ corresponds to an ideal $\mathscr{C}:=\eta^{-1}(\mathscr{C}_{1},\mathscr{C}_{2},\ldots,\mathscr{C}_{k})$ in $\mathscr{C}$ in $\mathscr{R}[x]/\langle f(x)\rangle$.
  2. 2.
    It follows directly from the proof of the first statement.
  3. 3.
    We apply the second statement in Proposition 3.
  4. 4.
    It follows from the proof of Proposition 3.
  5. 5.
    It is straightforward from the first statement.
  6. 6.
    Let $r_{i}=\operatorname{rank}_{\mathscr{R}_{i}}(\mathscr{C}_{i})$ and $w_{1}^{i},w_{2}^{i},\ldots,w_{r_{i}}^{i}$ be a set of generators for $\mathscr{C}_{i}$. Let $r=\max\{r_{1},r_{2},\ldots,r_{k}\}$. Consider the set of vectors
    $\left\{\eta^{-1}(v_{1}^{1},v_{1}^{2},\ldots,v_{1}^{k}),\eta^{-1}(v_{2}^{1},v_{2}^{2},\ldots,v_{2}^{k}),\ldots,\eta^{-1}(v_{r}^{1},v_{r}^{2},\ldots,v_{r}^{k})\right\}$
    where $v_{j}^{i}$ is the $j$-th component of $w_{j}^{i}$ and zeros are added as necessary to make each vector have $k$ components. This set spans $\mathscr{C}$. To see this, note that each generator $w_{j}^{i}$ can be mapped via $\eta$ to a vector in $\mathscr{C}$ by placing it in the $i$-th component and filling the other components with zeros. Thus, this set generates $\mathscr{C}$. Since we require $r$ vectors to generate $\mathscr{C}$ and there exists at least one index $i$ such that $r_{i}=r$, therefore it follows that $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})=\max\{\operatorname{rank}_{\mathscr{R}_{i}}(\mathscr{C}_{i})\mid i=1,\ldots,k\}$.
  7. 7.
    Let $d_{\star}=\min\{d(\mathscr{C}_{i})\mid i=1,\ldots,k\}$. Then, there exists a $j\in\{1,\ldots,k\}$ such that $d(\mathscr{C}_{j})=d_{\star}$. Let $v_{j}$ be a minimum weight non-zero vector in $\mathscr{C}_{j}$. Consider the vector $\eta^{-1}(0,0,\cdots,0,v_{j},0,\ldots,0)$, where $v_{j}$ is in the $j$-th component. This vector has Hamming weight $d_{\star}$, which implies $d(\mathscr{C})\leqslant d_{\star}$. Now, let $v$ be a minimum weight vector in $\mathscr{C}$. Then $\eta(v)$ has components $(v_{1},v_{2},\ldots,v_{k})$, where each $v_{i}$ belongs to $\mathscr{C}_{i}$. Since $d_{\star}$ is the minimum distance of $\mathscr{C}_{j}$, the weight of $v_{j}$ is at least $d_{\star}$. Thus, the weight of $\eta(v)$ is at least $d_{\star}$, which implies $d(\mathscr{C})\geqslant d_{\star}$. Therefore, $d_{\star}=d(\mathscr{C})$.
  8. 8.
[…]
4 Description of polycyclic codes over $\mathscr{R}$ as $R$-submodules
A monomial $\mathbf{x}^{\alpha}$ in $R[x_{1},x_{2},\ldots,x_{s}]$, where $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{s})\in(\mathbb{N}\cup\{0\})^{s}$, is a product of the form $\mathbf{x}^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\ldots x_{s}^{\alpha_{s}}$. We fix the lexicographic ordering on the monomials induced from the lexicographic ordering on ($(\mathbb{N}\cup\{0\})^{s}$, that is, $\mathbf{x}^{\alpha}{\succ}\mathbf{x}^{\beta}$ if the leftmost nonzero entry in the vector difference $\alpha-\beta$ is positive. Any element $g(x_{1},x_{2},\ldots,x_{s})\in R[x_{1},x_{2},\ldots,x_{s}]$ can be expressed uniquely as
$g(x_{1},x_{2},\ldots,x_{s})=\sum\limits_{\alpha\in(\mathbb{N}\cup\{0\})^{s}}g_{\alpha}\mathbf{x}^{\alpha},$ (8)
where finitely many coefficients $g_{\alpha}$ are non-zero. The degree of the monomial $\mathbf{x}^{\alpha}$ is $\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s}$ and the degree of $g(x_{1},x_{2},\ldots,x_{s})$ is the maximum degree among the degrees of the monomials involved in $g(x_{1},x_{2},\ldots,x_{s})$. Note that, fixed the lexicographic ordering, the sorted list of monomials can be represented as $\{\mathbf{x}^{\alpha^{(1)}},\mathbf{x}^{\alpha^{(2)}},\ldots,\mathbf{x}^{\alpha^{(m)}}\}$, where $m=m_{1}m_{2}\ldots m_{s}$ and $\alpha^{(j)}_{i}<m_{i}$ the degree of $t_{i}(x_{i})$ is a basis for $\mathscr{R}$. Therefore, we can represent an element $g(x_{1},x_{2},\ldots,x_{k})\in\mathscr{R}$ as $\sum_{i=1}^{m}g_{i}\mathbf{x}^{\alpha^{(i)}}$. We define the $R$-isomorphism $\Phi:\mathscr{R}\to R^{m}$, which maps $g(x_{1},x_{2},\ldots,x_{s})$ to the vector $(g_{1},g_{2},\ldots,g_{m})$. Therefore, $\Psi:=\Phi\circ\phi^{-1}:\mathscr{R}_{f}\to R^{mn}$ is an $R$-module isomorphism, where $\phi$ is defined in Section 3. Through this mapping, $f(x)$-polycyclic codes over $\mathscr{R}$ can be viewed as $R$-submodules of $R^{mn}$. Consider $\mathscr{C}$ an $f(x)$-polycyclic code, and let $c(x)=c_{0}+c_{1}x+c_{2}x^{2}+\cdots+c_{n-1}x^{n-1}$ be an element of $\mathscr{C}$. When $s=2$, we take the lexicographical order
$\{1,\,x_{1},\,\ldots\,,x_{1}^{m_{1}-1},\,x_{2},\,x_{1}x_{2},\,\ldots\,,x_{1}^{m_{1}-1}x_{2},\,\ldots\,,x_{2}^{m_{2}-1},\,x_{1}x_{2}^{m_{2}-1},\,\ldots\,,x_{1}^{m_{1}-1}x_{2}^{m_{2}-1}\}.$ (9)
Hence, $\Psi(c)$ is represented as the vector $(\Phi(c_{0}),\Phi(c_{1}),\ldots,\Phi(c_{n-1}))$, where
$\Phi(c_{i})=(c^{0,0}_{i},\ldots,c^{m_{1}-1,0}_{i},c^{0,1}_{i},\ldots,c^{m_{1}-1,1}_{i},\ldots,c^{0,m_{2}-1}_{i},\ldots,c^{m_{1}-1,m_{2}-1}_{i})$ (10)
and $c^{u,v}_{i}$ are the coefficients of $x_{1}^{u}x_{2}^{v}$ for $10\leqslant u\leqslant m_{1}-$ 1 and $10\leqslant v\leqslant m_{2}-$ 1. Each $\Phi(c_{i})$ can be viewed as the following array
$\begin{pmatrix}c^{0,0}_{i}&\cdots&c^{m_{1}-1,0}_{i}\\ c^{0,1}_{i}&\cdots&c^{m_{1}-1,1}_{i}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{i}&\cdots&c^{m_{1}-1,m_{2}-1}_{i}\\ \end{pmatrix}.$ (11)
Therefore, $\Psi(c)$ consist of $n$-subarrays arranged as follows:
($\begin{matrix}c^{0,0}_{0}&\cdots&c^{m_{1}-1,0}_{0}\\ c^{0,1}_{0}&\cdots&c^{m_{1}-1,1}_{0}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{0}&\cdots&c^{m_{1}-1,m_{2}-1}_{0}\\ \end{matrix}&\begin{matrix}c^{0,0}_{1}&\cdots&c^{m_{1}-1,0}_{1}\\ c^{0,1}_{1}&\cdots&c^{m_{1}-1,1}_{1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{1}&\cdots&c^{m_{1}-1,m_{2}-1}_{1}\\ \end{matrix}&\,\,\cdots&\begin{matrix}c^{0,0}_{n-1}&\cdots&c^{m_{1}-1,0}_{n-1}\\ c^{0,1}_{n-1}&\cdots&c^{m_{1}-1,1}_{n-1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{n-1}&\cdots&c^{m_{1}-1,m_{2}-1}_{n-1}\\ \end{matrix}\\ \end{array}\end{pmatrix}.$ (12)
Remark 2.
Since $\mathscr{C}$ is an $f(x)$-polycyclic code, it follows that $x_{1}c\in\mathscr{C}$, $x_{2}c\in\mathscr{C}$, and $xc\in\mathscr{C}$. The inclusion of $x_{1}c\in\mathscr{C}$ implies that the columns of each subarray in (12) are closed under $t_{1}(x_{1})$-polycyclic shift. Similarly, the inclusion of $x_{2}c$ in $\mathscr{C}$ implies that the rows of each subarray in Equation (12) are closed under $t_{2}(x_{2})$-polycyclic shift. Furthermore, the inclusion of $xc\in\mathscr{C}$ implies that the subarrays in Equation (12) themselves are closed under $f(x)$-polycyclic shift.
Definition 1.
Let $m=m_{1}m_{2}$. A length $mn$ linear code $C$ over the chain ring $R$ with codewords given as in Equation (12) is called quasi-2-dimensional polycyclic (Q2DP) code if the columns of each subarray are closed under $t_{1}(x_{1})$-polycyclic shift and the rows of each subarray are closed under $t_{2}(x_{2})$-polycyclic shift.
The above definition is a generalization of quasi-2-dimensional cyclic (Q2DC) codes defined in [18]. Similarly, we can define a multidimensional version of Q2DP codes and abbreviate them as QsDP.
Definition 2.
A Q2DP code is called an $f(x)$-polycyclic-Q2DP code if the subarrays in Equation (12) are closed under the $f(x)$-polycyclic shift.
Similarly, we can define a multidimensional version of $f(x)$-polycyclic-Q2DP codes, termed as $f(x)$-polycyclic-QsDP codes. Based on Remark 2, we have the following theorem.
Theorem 3.
Any $f(x)$-polycyclic code over $\mathscr{R}$ is $R$-isomorphic to an $f(x)$-polycyclic-QsDP code.
We know that the basis of the algebra $\mathscr{R}_{f}$ as an $\mathscr{R}$-module is $\{1,x,\ldots,x^{n-1}\}$. Recall that $\phi^{-1}$ maps the polynomial $g(x)=g_{0}+g_{1}x+\cdots+g_{n-1}x^{n-1}\in\mathscr{R}_{f}$ to $(g_{0},g_{1},\ldots,g_{n-1})$. The representation matrix of $g(x)$ is a matrix whose rows are $\phi^{-1}(g(x)),\phi^{-1}(xg(x)),\ldots,\phi^{-1}(x^{n-1}g(x))$. It is easy to see that the representation matrix of $x$ is the companion matrix $C_{f}$. So, similar to [5, 23], we can prove that any $f(x)$-polycyclic code over $\mathscr{R}$, as an $\mathscr{R}$-submodule of $\mathscr{R}^{n}$, is invariant under right multiplication by the companion matrix $C_{f}$. Now, our objective is to explore the relationship between $f(x)$-polycyclic codes (or equivalently, $f(x)$-polycyclic-QsDP codes) and invariant $R$-submodules of $R^{mn}$.
The basis of $\mathscr{R}_{f}$ as an $\mathscr{R}$-module is $\{1,x,\ldots,x^{n-1}\}$ whereas its basis as an $R$-module is the set
[…]
For example, suppose that $\mathscr{R}=R[x_{1},x_{2}]/\langle t_{1}(x_{1}),t_{2}(x_{2})\rangle$, where $t_{1}(x_{1})=x_{1}^{2}-b_{1}x_{1}-b_{0}$, $t_{2}(x_{2})=x_{2}^{2}-c_{1}x_{2}-c_{0}$ and $f(x)=x^{2}-f_{1}x-f_{0}$. If we take the lexicographic order $\{1,x_{1},x_{2},x_{1}x_{2}\}$ over $\mathscr{R}$, then the basis of $\mathscr{R}_{f}$ as an $R$-module is =$\mathcal{B}=\{1,x_{1},x_{2},x_{1}x_{2},x,x_{1}x,x_{2}x,x_{1}x_{2}x\}$. The representation matrix of $x$ with respect to $\mathcal{B}$ is a matrix whose rows consist of coefficients of polynomials $x,x_{1}x,x_{2}x,x_{1}x_{2}x,x^{2},x_{1}x^{2},x_{2}x^{2},x_{1}x_{2}x^{2}$, respectively. We denote this matrix by $C_{F}$ and let $\Phi(f_{i})=(r_{0}^{i},r_{1}^{i},r_{2}^{i},r_{3}^{i})$ for $i=0,1$. Then, it is easy to see that
$C_{F}=\begin{pmatrix}O_{4\times 4}&I_{4\times 4}\\ F_{0}&F_{1}\\ \end{pmatrix},$
where $O_{4\times 4}$ denotes the zero $4\times 4$ matrix, $I_{4\times 4}$ the identity $4\times 4$ matrix and
$F_{0}=\begin{pmatrix}{\Phi(f_{0})\bigg{(}I_{2\times 2}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(f_{0})\bigg{(}I_{2\times 2}\otimes C_{t_{1}}\bigg{)}}\\ {\Phi(f_{0})\bigg{(}C_{t_{2}}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(f_{0})\bigg{(}C_{t_{2}}\otimes C_{t_{1}}\bigg{)}}\\ \end{pmatrix},\quad F_{1}=\begin{pmatrix}{\Phi(f_{1})\bigg{(}I_{2\times 2}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(f_{1})\bigg{(}I_{2\times 2}\otimes C_{t_{1}}\bigg{)}}\\ {\Phi(f_{1})\bigg{(}C_{t_{2}}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(f_{1})\bigg{(}C_{t_{2}}\otimes C_{t_{1}}\bigg{)}}\\ \end{pmatrix}.$
Here $C_{t_{1}}$ and $C_{t_{2}}$ are companion matrices associated with $t_{1}(x_{1})$ and $t_{2}(x_{2})$, respectively. We notice that
$O_{4\times 4}=\begin{pmatrix}{\Phi(0)\bigg{(}I_{2\times 2}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(0)\bigg{(}I_{2\times 2}\otimes C_{t_{1}}\bigg{)}}\\ {\Phi(0)\bigg{(}C_{t_{2}}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(0)\bigg{(}C_{t_{2}}\otimes C_{t_{1}}\bigg{)}}\\ \end{pmatrix},\quad I_{4\times 4}=\begin{pmatrix}{\Phi(1)\bigg{(}I_{2\times 2}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(1)\bigg{(}I_{2\times 2}\otimes C_{t_{1}}\bigg{)}}\\ {\Phi(1)\bigg{(}C_{t_{2}}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(1)\bigg{(}C_{t_{2}}\otimes C_{t_{1}}\bigg{)}}\\ \end{pmatrix}.$
If we denote by $\Phi(C_{f})$ the matrix obtained by applying the map $\Phi$ to each entry of the companion matrix $C_{f}$, it becomes evident that rows 1 and 5 of $C_{F}$ form $\Phi(C_{f})\otimes\big{(}I_{2\times 2}\otimes I_{2\times 2}\big{)}$, rows 2 and 6 form $\Phi(C_{f})\otimes\big{(}I_{2\times 2}\otimes C_{t_{1}}\big{)}$, rows 3 and 7 form $\Phi(C_{f})\otimes\big{(}C_{t_{2}}\otimes I_{2\times 2}\big{)}$, and rows 4 and 8 form $\Phi(C_{f})\otimes\big{(}C_{t_{2}}\otimes C_{t_{1}}\big{)}$. Therefore, $C_{F}$ is equivalent to the following matrix:
$\begin{pmatrix}{\Phi(C_{f})\bigg{(}I_{2\times 2}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(C_{f})\bigg{(}I_{2\times 2}\otimes C_{t_{1}}\bigg{)}}\\ {\Phi(C_{f})\bigg{(}C_{t_{2}}\otimes I_{2\times 2}\bigg{)}}\\ {\Phi(C_{f})\bigg{(}C_{t_{2}}\otimes C_{t_{1}}\bigg{)}}\\ \end{pmatrix}=\Phi(C_{f})\otimes\begin{pmatrix}{I_{2\times 2}\otimes I_{2\times 2}}\\ {I_{2\times 2}\otimes C_{t_{1}}}\\ {C_{t_{2}}\otimes I_{2\times 2}}\\ {C_{t_{2}}\otimes C_{t_{1}}}\\ \end{pmatrix}=\Phi(C_{f})\otimes\begin{pmatrix}{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0}}\\ {C_{t_{2}}^{0}\otimes C_{t_{1}}}\\ {C_{t_{2}}\otimes C_{t_{1}}^{0}}\\ {C_{t_{2}}\otimes C_{t_{1}}}\\ \end{pmatrix}:=\Phi(C_{f})\otimes M.$ (14)
Thus, if we consider $\mathscr{R}_{f}$ as an $R$-module, then the representation matrix of $x$ is the matrix in Equation (14). Hence, Keeping the above notations, the following results hold for the $s=2$ case.
Theorem 4.
Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$. Then $\Psi(\mathscr{C})$ is closed under the right multiplication by the matrix $\Phi(C_{f})\otimes M$, defined in (14).
Corollary 1.
Any $f(x)$-polycyclic-Q2DP code is closed under the right multiplication by $\Phi(C_{F})\otimes M$.
Similarly, the above theorem and corollary are true for any value of $s$. It is worth noting that $\{1,x_{1},x_{2},x_{1}x_{2}\}$ forms the basis for $\mathscr{R}$, and the representation matrices for these basis elements are $\{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0},C_{t_{2}}^{0}\otimes C_{t_{1}},C_{t_{2}}\otimes C_{t_{1}}^{0},C_{t_{2}}\otimes C_{t_{1}}\}$, see [5, Section 6]. We notice that the matrix $M$ is constructed based on this set. Therefore, for the general case $s$, the matrix $M$ can be constructed using the representation matrices of the basis elements of $\mathscr{R}$ and $R$-module.
5 Euclidean dual of polycyclic codes over $\mathscr{R}$
A linear code $\mathscr{C}$ of length $n$ over $\mathscr{R}$ is called $f(x)$-sequential if it is invariant under the $f(x)$-sequential shift $\nu_{f(x)}:\mathscr{R}^{n}\rightarrow\mathscr{R}^{n}$ defined by
$\nu_{f(x)}(b_{0},b_{1},\ldots,b_{n-1})=(b_{1},b_{2},\ldots,b_{n-2},d),$ (15)
where $d=f_{0}b_{0}+f_{1}b_{1}+\cdots+f_{n-1}b_{n-1}$. As usual, the Euclidean product of $x$ and $y$ in $\mathscr{R}^{n}$ is given by $\langle x,y\rangle=x\cdot y^{\text{tr}}$, where $y^{\text{tr}}$ denotes the transpose of $y$. The Euclidean dual code $\mathscr{C}^{\perp}$ of an $\mathscr{R}$-linear code $\mathscr{C}$ of length $n$ is defined as $\mathscr{C}^{\perp}:=\{x\in\mathscr{R}^{n}\mid\langle x,y\rangle=0,\,\forall y\in\mathscr{C}\}$. It is easy to see that the operator $\nu_{f(x)}$ is the adjoint operator of $\tau_{f(x)}$, that is $\langle\tau_{f(x)}(x),y\rangle=\langle x,\nu_{f(x)}(y)\rangle$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator Therefore, we have the following theorem.
Theorem 5.
The linear code $\mathscr{C}$ is an $f(x)$-polycyclic code over $\mathscr{R}$ if and only if $\mathscr{C}^{\perp}$ is an $f(x)$-sequential code.
Theorem 6.
[23, Theorem 3.5] The codes $\mathscr{C}$ and $\mathscr{C}^{\perp}$ are $f(x)$-polycyclic codes if and only if they are both $f(x)$-sequential codes, if and only if $\mathscr{C}$ is a constacyclic code.
Using the results in [24], we can establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case of $f(x)$. Recall that $R$ is a chain ring with the maximal ideal $\gamma$ and nilpotency index $e$, $\mathcal{C}$ is the partition of $\mathcal{H}=\prod_{i=1}^{s}H_{i}$, and $h_{C}$ is defined in Equation (4). We denote the ideal $\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle\subseteq R[x_{1},\ldots,x_{s}]$ by $I$.
Theorem 7.
Let $f(x)$ be a regular basic irreducible polynomial with coefficients in $R$. An $f(x)$-polycyclic code over the ring $\mathscr{R}$ is isomorphic to a sum of ideals of $R_{f}[x_{1},\ldots,x_{s}]/I$ of the form
$\left\langle\big{(}\gamma+\langle f(x)\rangle\big{)}^{j_{C}}h_{C}(x_{1},x_{2},\ldots,x_{s})+I\right\rangle,$
where $0\leqslant j_{C}\leqslant e$ and $R_{f}=R[x]/\langle f(x)\rangle$.
Proof.
If $f(x)$ is a regular basic irreducible polynomial in the ring $R[x]$, then $R_{f}:=R/\langle f(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma+\langle f(x)\rangle\rangle$, see Remark 1. It is well-known that
$\frac{\mathscr{R}[x]}{\langle f(x)\rangle}\cong\frac{\frac{R[x]}{\langle f(x)\rangle}[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}=\frac{R_{f}[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}.$ (16)
Hence, every $f(x)$-polycyclic code over $\mathscr{R}$ is isomorphic to an ideal in the multivariable polynomial ring $R_{f}[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$ over the chain ring $R_{f}$. Now use Corollary 1 in [24]. ∎
Remark 3.
If $f(x)$ is a monic square-free polynomial with coefficients in $R$, it can be factored as a product of regular basic irreducible pairwise coprime polynomials with coefficients in $R$. Therefore, by the Chinese Remainder Theorem, we can extend the above theorem for a monic square-free polynomial with coefficients in $R$.
Theorem 8.
Let $f(x)$ be a regular basic irreducible polynomial with coefficients in $R$. Then the following statements hold.
  1. 1.
    The number of $f(x)$-polycyclic codes over $\mathscr{R}$ is equal to $(e+1)^{k}$, where $k=\lvert\mathcal{C}\rvert$.
  2. 2.
    For the $f(x)$-polycyclic code $\mathscr{C}$, there exists a family of polynomials $G_{0},\ldots,G_{s}\in R[x_{1},\ldots,x_{s}]$ such that
    $\mathscr{C}\cong\left\langle G_{1},\,\,\big{(}\gamma+\langle f(x)\rangle\big{)}G_{2},\,\,\ldots\,\,,\big{(}\gamma+\langle f(x)\rangle\big{)}^{e-1}G_{s}\right\rangle+I.$ (17)
  3. 3.
    The size of the $f(x)$-polycyclic code $\mathscr{C}$ is equal to
    $\bigg{|}\frac{R_{f}}{\langle\gamma+\langle f(x)\rangle\rangle}\Bigg{|}^{\sum_{i=0}^{e-1}(e-i)N_{i}},$
    where $N_{i}$ denotes the number of zeros $\mu\in\prod_{i=1}^{s}H_{i}$ of $\overline{G}_{i}$. Here, $\overline{G}_{i}$ represents the natural ring homomorphism image over $R_{f}/\langle\gamma+\langle f(x)\rangle\rangle$.
Proof.
Based on the proof of Theorem 7 , every $f(x)$-polycyclic code over $\mathscr{R}$ is isomorphic to an ideal in the multivariable polynomial ring $R_{f}[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$ over the chain ring $R_{f}$. Now, we apply Corollary 2, Theorem 3 and Corollary 3 in [24]. ∎
Theorem 9 ($\mathscr{R}$ Abelian).
Let $f(x)$ be a regular basic irreducible polynomial with coefficients in $R$, and let $t_{i}(x_{i})$ be square-free polynomials of the form $t_{i}(x_{i})=x_{i}^{e_{i}}-1$ for $1\leqslant i\leqslant s$, and the ring $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$. Then the following statements hold.
  1. 1.
    If $f(x)$-polycyclic code $\mathscr{C}$ has the form in Equation (17), then
    $\mathscr{C}^{\perp}=\left\langle\tau(G_{0}),\,\,\big{(}\gamma+\langle f(x)\rangle\big{)}\tau(G_{2}),\,\,\ldots\,\,,\big{(}\gamma+\langle f(x)\rangle\big{)}^{e-1}\tau(G_{s})\right\rangle+I,$ (18)
    where $\tau$ is the ring automorphism of $R_{f}[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$ given by $\tau(g(x_{1},x_{2},\ldots,x_{s}))=g(x_{1}^{e_{1}-1},x_{1}^{e_{2}-1},\ldots,x_{1}^{e_{s}-1})$.
  2. 2.
    The size of $\mathscr{C}^{\perp}$ is equal to
    $\bigg{|}\frac{R_{f}}{\langle\gamma+\langle f(x)\rangle\rangle}\bigg{|}^{\sum_{i=0}^{e-1}iN_{i}},$
    where $N_{i}$ denotes the number of zeros $\mu\in\prod_{i=1}^{s}H_{i}$ of $\overline{G}_{i}$.
Proof.
See Theorem 4, Corollary 5 in [24]. ∎
6 Annihilator dual of polycyclic codes over $\mathscr{R}$
As pointed before, the Euclidean dual of an $f(x)$-polycyclic code over $\mathscr{R}$ is not necessarily an $f(x)$-polycyclic code. In [1], an alternative duality for $f(x)$-polycyclic codes over fields is introduced, where the dual code of any $f(x)$-polycyclic code preserves the $f(x)$-polycyclic structure. In this section, we consider this duality for $f(x)$-polycyclic codes over $\mathscr{R}$. Let $A$ be a commutative ring with $1$ and $k(x)$, $l(x)$ be two non-zero polynomials in $A[x]$ such that the leading coefficient of $l(x)$ is a unit in $A$. Then the division algorithm holds, that is, there exist unique polynomials $q(x)$, $r(x)\in A[x]$ such that $k(x)=q(x)l(x)+r(x)$ and either $r(x)=0$ or $\deg r(x)<\deg l(x)$.
Definition 3.
Let $\mathscr{C}$ be an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$. The annihilator product of $g(x),h(x)\in\mathscr{R}_{f}$ is defined as
$\langle g(x),h(x)\rangle_{f}=r(0),$
where $r(x)$ is the remainder of $g(x)h(x)$ divided by $f(x)$. Moreover, the annihilator dual of $\mathscr{C}$ is defined by $\mathscr{C}^{\circ}=\{h(x)\in\mathscr{R}_{f}\mid\langle g(x),h(x)\rangle_{f}=0,\forall g(x)\in\mathscr{C}\}.$
We notice that the definition is well-defined due to the uniqueness of the reminder in the division algorithm. The $f(x)$-polycyclic code $\mathscr{C}$ is called annihilator self-orthogonal (respectively annihilator self-dual, annihilator linear complementary dual (LCD), or annihilator dual-containing) if $\mathscr{C}\subseteq\mathscr{C}^{\circ}$ (respectively $\mathscr{C}=\mathscr{C}^{\circ}$, $\mathscr{C}\cap\mathscr{C}^{\circ}=\{0\}$, or $\mathscr{C}^{\circ}\subseteq\mathscr{C}$). Recall that
$\operatorname{Ann}(\mathscr{C})=\{h(x)\in\mathscr{R}_{f}\mid g(x)h(x)=0\,(\operatorname{mod}f(x)),\,\,\forall g(x)\in\mathscr{C}\}.$
Remark 4.
If $g(x)$ and $h(x)$ are two polynomials with coefficients in a ring $A$, then $\deg(g(x)h(x))\leqslant\deg(g(x))+\deg(h(x))$. If $\operatorname{lc}(g)\operatorname{lc}(h)\neq 0$, then the equality holds, where “$\operatorname{lc}$” denotes the leading coefficient of each polynomial.
Theorem 10.
Let $f_{0}$ be invertible in $\mathscr{R}$ and $\mathscr{C}$ be an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$.
  1. 1.
    The annihilator product is a non-degenerate $\mathscr{R}$-bilinear form.
  2. 2.
    The annihilator dual $\mathscr{C}^{\circ}$ is equal to $\operatorname{Ann}(\mathscr{C})$.
  3. 3.
    Let the leading coefficient of $g(x)\in\mathscr{R}_{f}$ be a unit element and $\mathscr{C}$ be a $f(x)$-polycyclic code generated by $g(x)$. If there exists a polynomial $h(x)\in\mathscr{R}_{f}$ satisfying $f(x)=g(x)h(x)$, then $\mathscr{C}^{\circ}=\langle h(x)\rangle$. Moreover, $|\mathscr{C}|\,|\mathscr{C}^{\circ}|=|\mathscr{R}|^{n}$.
Proof.
  1. 1.
    It is straightforward that the given product is an $\mathscr{R}$-bilinear form. Moreover, it is easy to see that $x$ is invertible with inverse $m(x)=\sum_{i=0}^{n-2}-f_{0}^{-1}f_{i+1}x^{i}+f_{0}^{-1}x^{n-1}.$ To prove the non-degeneracy, suppose that $g(x)=g_{0}+g_{1}x+\cdots+g_{n-1}x^{n-1}\in\mathscr{R}_{f}$ satisfies $\langle g(x),h(x)\rangle_{f}=0$ for all $h(x)\in\mathscr{R}_{f}.$ In the case $h=1$, we have $0=r(0)=g_{0}$. Hence, we can rewrite $g(x)$ as $(g_{1}+g_{2}x+\cdots+g_{n-1}x^{n-2})x$. In the case $h=x^{-1}$, we get $g_{1}=0$. Repeating this discussion, we conclude that $g_{i}=0$ for all $i$, which implies $g(x)=0$. Therefore, the annihilator product is non-degenerate.
  2. 2.
    It is easy to see that $\operatorname{Ann}(\mathscr{C})\subseteq\mathscr{C}^{\circ}.$ Conversely, let $g(x)\in\mathscr{C}^{\circ}$ and $h(x)$ be an arbitrary element of $\mathscr{C}$. Let $r(x)=r_{0}+r_{1}x+\cdots+r_{n-1}x^{n-1}$ denote the remainder of $g(x)h(x)$ modulo $f(x)$. We have $\langle g(x),h(x)\rangle_{f}=0$, which implies $r_{0}=0$. Consequently, $(r_{1}+r_{2}x+\cdots+r_{n-1}x^{n-2})x=g(x)h(x)(\operatorname{mod}f(x)).$ According to the proof of the first statement, the inverse of $x$ is $m(x)$. Thus, $r_{1}+r_{2}x+\cdots+r_{n-1}x^{n-2}=g(x)h(x)m(x)(\operatorname{mod}f(x))$. Furthermore, $ghm(0)=0(\operatorname{mod}f(x))$, which implies $r_{1}=0$. By repeating the same reasoning, we have $r_{i}=0$ for all $i$, thus $gh=0$.
  3. 3.
    Since $f(x)=g(x)h(x)$, $\langle h(x)\rangle\subseteq\operatorname{Ann}(\langle g(x)\rangle)=\mathscr{C}^{\circ}$. Conversly, let $k(x)\in\operatorname{Ann}(\langle g(x)\rangle)$, then there exists a polynomial $t(x)$ such that $k(x)g(x)=t(x)f(x)$. If $\operatorname{lc}(g)\operatorname{lc}(h)=0$, then $\operatorname{lc}(h)=0$, which is a contradiction. Therefore, $\operatorname{lc}(g)\operatorname{lc}(h)\neq 0$, which implies that $\deg(g(x)h(x))=\deg(g(x))+\deg(h(x))$. Consequently, since $g(x)h(x)=f(x)$, we deduct $\operatorname{lc}(g)\operatorname{lc}(h)=\operatorname{lc}(f)=1$, which implies $\operatorname{lc}(h)$ is invertible. Therefore, we can apply the division algorithm to the monic polynomial $\operatorname{lc}(h)^{-1}h(x)$, that is, there exist unique elements $q(x),r(x)\in\mathscr{R}_{f}$ such that $k(x)=q(x)\operatorname{lc}(h)^{-1}h(x)+r(x)$ and $\deg(r(x))<\deg(\operatorname{lc}(h)^{-1}h(x))$. So, $k(x)g(x)=q(x)a^{-1}f(x)+r(x)g(x)$. This leads to $(t(x)-q(x)a^{-1})f(x)=r(x)g(x)$, and hence $r(x)g(x)\in\langle f(x)\rangle$ Assume that $r(x)$ is non-zero. As $\operatorname{lc}(g)$ is unit, it follows that $r(x)g(x)\neq 0$. Consequently, $0\neq r(x)g(x)\in\langle f(x)\rangle$, which contradicts the fact that $h(x)$ is the smallest polynomial satisfying $g(x)h(x)=f(x)$ and $\deg(r(x))<\deg(\operatorname{lc}(h)^{-1}h(x))=\deg(h(x))$. Thus, $r(x)=0$, and hence $k(x)\in\langle h(x)\rangle$. As a result, $\operatorname{Ann}(\langle g(x)\rangle)\subseteq\langle h(x)\rangle$. To complete the proof, note that $\langle f(x)/g(x)\rangle\oplus\langle g(x)\rangle=\mathscr{R}[x]/\langle f(x)\rangle$. Therefore, $|\mathscr{C}||\mathscr{C}^{\circ}|=|\mathscr{R}|^{n}$.
The set of vectors $\{e_{i}:1\leqslant i\leqslant n\}$, where each $e_{i}$ has a value of $1$ in the $i$-th position and 00 elsewhere, forms a basis for $\mathscr{R}^{n}$. The Gram matrix $A$ associated with this set is a matrix with entries $a_{i,j}=\langle e_{i},e_{j}\rangle_{f}$. For a linear code $\mathscr{C}$ over $\mathscr{R}$, we define the set $\mathscr{C}\cdot A:=\{cA\mid c\in\mathscr{C}\}$. Similar to Lemma 4 and Corollary 1 in [14], we have the following Lemma.
Lemma 2.
Let $f_{0}\in\mathscr{R}$ be invertible element and $\mathscr{C}$ be an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$.
  1. 1.
    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$.
  2. 2.
    $(\mathscr{C}^{\circ})^{\circ}=\mathscr{C}$.
Corollary 2.
Let $f_{0}\in\mathscr{R}$ be an invertible element and $\mathscr{C}$ be a linear code of length $n$ over $\mathscr{R}$. Then, $\mathscr{C}$ is an $f(x)$-polycyclic code if and only if $\mathscr{C}^{\circ}$ is an $f(x)$-polycyclic code.
Proof.
It follows from the fact that $\operatorname{Ann}(\mathscr{C})$ is an ideal of $\mathscr{R}_{f}$, and $(\mathscr{C}^{\circ})^{\circ}=\mathscr{C}$. ∎
Theorem 11.
Let $f_{0}\in\mathscr{R}$ be an invertible element and $\mathscr{C}$ be an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ with the decomposition $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}$.
  1. 1.
    $\mathscr{C}^{\circ}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}$.
  2. 2.
    $\mathscr{C}$ is annihilator self-orthogonal (resp. self-dual, LCD, dual-containing) if and only if each $\mathscr{C}_{i}$ is annihilator self-orthogonal (resp. self-dual, LCD, dual-containing).
Proof.
  1. 1.
    First, we claim that the annihilator product defined on each $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is non-degenerate. To prove it, note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the finite ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$, see [28, Chapter 1]. Thus, since the constant part of the polynomial $f(x)$ is invertible, the constant part of each polynomial $f_{i}(x)$ is also invertible, and therefore, by Theorem 10 the claim follows. According to the proof of the first statement in Proposition 3, the ring $\mathscr{R}[x]/\langle f(x)\rangle\cong\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is isomorphic to $\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle e_{C_{i}}$, thus $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}$. Furthermore, since the entries of the Gram matrix $A$ are in $\mathscr{R}$, by Proposition 1 yields $A=\oplus_{i=1}^{k}A_{i}e_{C_{i}}$, where $A_{i}$ is an $n\times n$ matrices over $\mathscr{R}$. Consequently, by Lemma 2, we have
    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$ $=\bigg{(}\big{(}\bigoplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\big{)}\cdot\big{(}\bigoplus_{i=1}^{k}A_{i}e_{C_{i}}\big{)}\bigg{)}^{\perp}$
    $=\bigg{(}\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}e_{C_{i}}\bigg{)}^{\perp}$
    $=\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}^{\perp}e_{C_{i}}=\bigoplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}.$
  2. 2.
    If $\mathscr{C}$ is annihilator self-orthogonal, then $\oplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\subseteq\oplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}$. Multiplying the idempotent orthogonal $e_{C_{i}}$ on both sides implies $\mathscr{C}_{i}\subseteq\mathscr{C}_{i}^{\circ}$ for all $i$. The proof of the other statements follows a similar reasoning.
Theorem 12.
Suppose Let $f_{0}\in\mathscr{R}$ be an invertible element, $\mathscr{C}$ is an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ generated by a single polynomial $g(x)\in\mathscr{R}_{f}$, and the leading coefficient of $g(x)$ be a unit element in $\mathscr{R}$ and there exist $h(x)\in\mathscr{R}_{f}$ such that $f(x)=g(x)h(x)$. Then
  1. 1.
    $\mathscr{C}$ is annihilator self-dual if and only if $f(x)=ag^{2}(x)$ for some unit element $a$ in $\mathscr{R}$.
  2. 2.
    $\mathscr{C}$ is annihilator self-orthogonal (dual-containing) if and only if $h(x)\mid g(x)$ ($g(x)\mid h(x)$).
  3. 3.
    $\mathscr{C}$ is annihilator LCD if and only if $\operatorname{gcd}(g(x),h(x))=1$.
Proof.
  1. 1.
    If $f(x)=ag^{2}(x)$, according to Theorem 10, we have $\mathscr{C}^{\circ}=\langle ag(x)\rangle=\langle g(x)\rangle=\mathscr{C}$. Conversely, let $\mathscr{C}$ be annihilator self-dual. Then, $g(x)\in\mathscr{C}$ implies $g(x)\in\mathscr{C}^{\circ}$. Thus, $g^{2}(x)=0(\operatorname{mod}f(x))$, that means there exists $l(x)$ in $\mathscr{R}[x]$ satisfying
    $g^{2}(x)=l(x)f(x).$ (19)
    On the other side, by Theorem 10, we have $\mathscr{C}^{\circ}=\langle h(x)\rangle$. Since $\mathscr{C}^{\circ}=\mathscr{C}$, it follows that $\langle h(x)\rangle=\langle g(x)\rangle$. Thus, there exists $k(x)$ in $\mathscr{R}[x]$ such that $h(x)=k(x)g(x)$. Consequently,
    $f(x)=k(x)g^{2}(x).$ (20)
    Now, since $\operatorname{lc}(g)\operatorname{lc}(g)\neq 0$, $\operatorname{lc}(l)\operatorname{lc}(f)\neq 0$ and $\operatorname{lc}(k)\operatorname{lc}(g^{2})\neq 0$, employing Remark 4 and Eqs. (19) and (20), we deduce $2\deg g=\deg l+\deg f$ and $\deg f=\deg k+2\deg g$, which imply $\deg l+\deg k=0$. Consequently, $\deg l=\deg k=0$. Thus, there exist elements $a,b\in R$ such that $l(x)=a$ , $k(x)=b$. Note that since leading coefficients of $f,g$ are units, Eq.s (19) and (20) yield $a$ and $b$ are units.
  2. 2.
    Using Theorem 10, we obtain $\mathscr{C}=\langle g(x)\rangle$ and $\mathscr{C}^{\circ}=\langle h(x)\rangle$, which complete the proof.
  3. 3.
    Similar to the proof of Theorem 2.8 in [32].
Remark 5.
In Theorem 3 of [1], it is asserted “A polycyclic code $\langle g\rangle$ in $\mathbb{F}[x]/\langle f\rangle$ is annihilator self-dual if and only if $f=g^{2}$”. The authors provided a proof for only one direction: “If $f=g^{2}$, then the polycyclic code $\langle g\rangle$ in $\mathbb{F}[x]/\langle f\rangle$ is annihilator self-dual”. However, they omitted the proof for the converse statement. The proof of Theorem 12 in this paper shows that the converse statement of their theorem is not true in general. Similarly, in Theorem 2.7 of [32], the author attempted to prove Theorem 3 of [1] for the case $\mathbb{F}$ is replaced with the ring $A=\mathbb{F}_{q}+\mathbb{F}_{q}$ with $u^{2}=0$. The author’s proof relies on this statement: “If $h(x)$ divides $g(x)$ and $g(x)$ divides $h(x)$, then $h(x)$ equals $g(x)$”, without considering that this divisibility relation cannot be true for elements in the ring of polynomials in general.
Corollary 3.
With the conditions in Theorem 12. Suppose that $\eta(g(x))=\big{(}g_{1}(x),g_{2}(x),\ldots,g_{k}(x)\big{)}$ and $\eta(h(x))=\big{(}h_{1}(x),h_{2}(x),\ldots,h_{k}(x)\big{)}$, where the mapping $\eta$ defined in Equation (7).
  1. 1.
    $\mathscr{C}$ is an annihilator self-dual code if and only if $f_{i}(x)=a_{i}g_{i}^{2}(x)$, for some unit elements $a_{i}$ in $\mathscr{R}_{i}$, for all $i\in\{1,\ldots,k\}$.
  2. 2.
    $\mathscr{C}$ is an annihilator self-orthogonal (dual-containing) code if and only if $h_{i}(x)\mid g_{i}(x)$ ($g_{i}(x)\mid h_{i}(x)$) for all $i\in\{1,\ldots,k\}$.
  3. 3.
    $\mathscr{C}$ is an annihilator LCD code if and only if $\operatorname{gcd}(g_{i}(x),h_{i}(x))=1$ for all $i\in\{1,\ldots,k\}$.
Proof.
We will proof the first statement, and the other ones follow similarly. Note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$ [28, Chapter 1]. Thus, if the leading coefficient of $g(x)$ is a unit element in $\mathscr{R}$, then the leading coefficient of $g_{i}(x)$ is a unit element in $\mathscr{R}_{i}$ for all $i\in\{1,\ldots,k\}$. Therefore, by Theorem 12, each $\mathscr{C}_{i}$ is an annihilator self-dual code if and only if $f_{i}(x)=a_{i}g_{i}^{2}(x)$ for some unit element $a_{i}$ in $\mathscr{R}_{i}$. Then, the proof follows by Theorem 11. ∎
Example 6.1.
Consider the ring $\mathscr{R}=R[x_{1},x_{2},x_{3}]/\langle x_{1}^{2}-1,x_{2}^{2}-1,x_{3}^{2}-1\rangle$, where $R$ is a chain ring with the residue field $\mathbb{F}_{q}$. Any element $r\in\mathscr{R}$ of the form $r=a_{1}+a_{2}x_{1}+a_{3}x_{2}+a_{4}x_{3}+a_{5}x_{1}x_{2}+a_{6}x_{2}x_{3}+a_{7}x_{1}x_{3}+a_{8}x_{1}x_{2}x_{3}$ can be uniquely expressed as $r_{1}e_{1}+r_{2}e_{2}+r_{3}e_{3}+r_{4}e_{4}+r_{5}e_{5}+r_{6}e_{6}+r_{7}e_{7}+e_{8}e_{8}$, where $r_{i}\in\mathscr{R}$ and
$e_{1}$ $=t(1+x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}+x_{1}x_{2}x_{3}),$
$e_{2}$ $=t(1+x_{1}+x_{2}-x_{3}+x_{1}x_{2}-x_{2}x_{3}-x_{1}x_{3}-x_{1}x_{2}x_{3}),$
$e_{3}$ $=t(1+x_{1}-x_{2}+x_{3}-x_{1}x_{2}-x_{2}x_{3}+x_{1}x_{3}-x_{1}x_{2}x_{3}),$
$e_{4}$ $=t(1-x_{1}+x_{2}+x_{3}-x_{1}x_{2}+x_{2}x_{3}-x_{1}x_{3}-x_{1}x_{2}x_{3}),$
$e_{5}$ $=t(1+x_{1}-x_{2}-x_{3}-x_{1}x_{2}+x_{2}x_{3}-x_{1}x_{3}+x_{1}x_{2}x_{3}),$
$e_{6}$ $=t(1-x_{1}-x_{2}+x_{3}+x_{1}x_{2}-x_{2}x_{3}-x_{1}x_{3}+x_{1}x_{2}x_{3}),$
$e_{7}$ $=t(1-x_{1}+x_{2}-x_{3}-x_{1}x_{2}-x_{2}x_{3}+x_{1}x_{3}+x_{1}x_{2}x_{3}),$
$e_{8}$ $=t(1-x_{1}-x_{2}-x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}-x_{1}x_{2}x_{3}),$
where $t\in\mathbb{F}_{q}$ and $8t=1\mod p$. As special case, let $R=\mathbb{Z}_{9}$ and $t=2$.
  1. 1.
    Suppose that $f_{0}=2x_{1}-2x_{2}-2x_{3}+x_{1}x_{2}x_{3}$, $f_{1}=x_{1}+2x_{3}$,$f_{2}=1-2x_{1}+2x_{2}+2x_{3}+x_{1}x_{2}+x_{2}x_{3}$, $g_{0}=x_{1}-2x_{3}+x_{1}x_{2}$, $g_{1}=-2x_{1}x_{2}-2x_{2}x_{3}-2x_{1}x_{3}$ and $g_{2}=1+x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}+x_{1}x_{2}x_{3}$ are elements in $\mathscr{R}$. Let $\mathscr{C}$ be a code in $\mathscr{R}[x]/\langle f(x)\rangle$ generated by $g(x)=g_{2}x^{2}+g_{1}x+g_{0}$, where $f(x)=x^{3}-f_{2}x^{2}-f_{1}x-f_{0}$. It is easy to see that
    $f_{0}=$ $2e_{1}+e_{2}+e_{5}+2e_{8},$ $g_{0}=$ $2e_{1}+2e_{2}+e_{5}+2e_{8},$
    $f_{1}=$ $2e_{4}+e_{5}+2e_{7}+e_{8},$ $g_{1}=$ $e_{1}+2e_{2}+2e_{3}+e_{4}+e_{6}+2e_{7},$
    $f_{2}=$ $2e_{4}+2e_{5}+e_{6}+e_{7}+e_{8},$ $g_{2}=$ $2e_{1}+2e_{2}+2e_{3}++2e_{4}+2e_{5}+2e_{6}+2e_{7}+2e_{8}.$
    Therefore,
    $f_{1}(x)$ $=x^{3}+1,$ $f_{2}(x)$ $=x^{3}+2,$ $f_{3}(x)$ $=x^{3}+x^{2},$ $f_{4}(x)$ $=x^{3}+x^{2}+x,$
    $f_{5}(x)$ $=x^{3}+x^{2}+2x+2,$ $f_{6}(x)$ $=x^{3}+2x^{2},$ $f_{7}(x)$ $=x^{3}+2x^{2}+x,$ $f_{8}(x)$ $=x^{3}+2x^{2}+2x+1,$
    and
    $g_{1}(x)$ $=2x^{2}+x+2,$ $g_{2}(x)$ $=2x^{2}+2x+2,$ $g_{3}(x)$ $=2x^{2}+2x,$ $g_{4}(x)$ $=2x^{2}+x,$
    $g_{5}(x)$ $=2x^{2}+1,$ $g_{6}(x)$ $=2x^{2}+x,$ $g_{7}(x)$ $=2x^{2}+2x,$ $g_{8}(x)$ $=2x^{2}+1.$
    One can easily see that $g^{2}_{i}(x)=f_{i}(x)$ for all $i$. Using Corollary 3, $\mathscr{C}$ is an annihilator self-dual code in $\mathscr{R}[x]/\langle f(x)\rangle$.
  2. 2.
    Suppose now that $f_{0}=-x_{3}-x_{1}x_{2}+x_{2}x_{3}-x_{1}x_{2}x_{3}$, $f_{1}=2x_{1}+x_{1}x_{2}+2x_{2}x_{3}+2x_{1}x_{3}+2x_{1}x_{2}x_{3}$, $f_{2}=1+2x_{2}-2x_{3}+2x_{1}x_{2}-2x_{1}x_{3}$, $g_{0}=2x_{1}+2x_{2}+x_{3}+x_{2}x_{3}+x_{1}x_{3}$, $g_{1}=2+x_{1}+2x_{2}+x_{3}+2x_{1}x_{2}+2x_{1}x_{3}$ and $g_{2}=2+x_{1}+2x_{2}+x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}+x_{1}x_{2}x_{3}$ are elements in $\mathscr{R}$. Let $\mathscr{C}$ be a code in $\mathscr{R}[x]/\langle f(x)\rangle$ generated by $g(x)=g_{2}x^{2}+g_{1}x+g_{0}$, where $f(x)=x^{3}-f_{2}x^{2}-f_{1}x-f_{0}$. It is easy to see that
    $f_{0}=$ $2e_{1}+2e_{2}+2e_{3}+2e_{4}+2e_{5}+2e_{6}+e_{7}+2e_{8},$ $g_{0}=$ $2e_{1}+2e_{2}+e_{5}+e_{8},$
    $f_{1}=$ $e_{1}+2e_{4}+2e_{6}+e_{8},$ $g_{1}=$ $2e_{1}+e_{4}+e_{6}+2e_{7},$
    $f_{2}=$ $2e_{2}+e_{3}+e_{4}+e_{8},$ $g_{2}=$ $2e_{1}+2e_{2}+2e_{3}++2e_{4}+2e_{5}+2e_{6}+2e_{7}+2e_{8}.$
    Therefore,
    $f_{1}(x)$ $=x^{3}+2x+1,$ $f_{2}(x)$ $=x^{3}+x^{2}+1,$ $f_{3}(x)$ $=x^{3}+2x^{2}+1,$ $f_{4}(x)$ $=x^{3}+2x^{2}+x+1,$
    $f_{5}(x)$ $=x^{3}+1,$ $f_{6}(x)$ $=x^{3}+x+1,$ $f_{7}(x)$ $=x^{3}+2,$ $f_{8}(x)$ $=x^{3}+2x^{2}+2x+1,$
    and
    $g_{1}(x)$ $=x^{2}+2x+2,$ $g_{2}(x)$ $=x^{2}+2x+2,$ $g_{3}(x)$ $=x^{2}+x+2,$ $g_{4}(x)$ $=x^{2}+x+2,$
    $g_{5}(x)$ $=x^{2}+2x+2,$ $g_{6}(x)$ $=x^{2}+x+2,$ $g_{7}(x)$ $=x^{2}+x+2,$ $g_{8}(x)$ $=x^{2}+2x+2.$
    Assume that $h_{i}(x)=x+1$. The greatest common divisor of $g_{i}(x)$ and $h_{i}(x)$ is $1$ for all $i$. Moreover, $f_{i}(x)=g_{i}(x)h_{i}(x)$. By Corollary 3, the code $\mathscr{C}$ generated by $g(x)$ is an annihilator LCD code in $\mathscr{R}[x]/\langle f(x)\rangle$.
7 Annihilator CSS construction
The CSS construction is a well-known method to construct quantum codes from linear codes over fields or rings, see for example [17, 29, 33]. Depending on the duality defined for the code, this technique is referred to as either the Euclidean CSS construction or the Hermitian CSS construction. As we noticed above, polycyclic codes are also equipped with the annihilator dual. In this section, we establish annihilator CSS construction, a novel approach for constructing quantum codes from annihilator dual-preserving polycyclic codes over finite chain rings.
Recall that since $R$ is a finite commutative chain ring with identity it is a Frobenius ring, thus it has generating characters. For the character $\chi$ in $\operatorname{Hom}(R;\mathbb{C}^{*})$, there exists a unique function $\psi$ in $\operatorname{Hom}(R;\mathbb{Q}/\mathbb{Z})$ such that $\chi(x)=\exp(2\pi i\psi(x))$. Let $\langle\,,\,\rangle_{\chi}:{R}^{2n}\times{R}^{2n}\rightarrow\mathbb{Q}/\mathbb{Z}$ and $\langle\,,\,\rangle_{s}:{R}^{2n}\times{R}^{2n}\rightarrow{R}$ be two forms defined as follows:
$\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{\chi}=\psi(b\cdot a^{\prime}-b^{\prime}\cdot a)\quad\text{and}\quad\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{s}=b\cdot a^{\prime}-b^{\prime}\cdot a.$
We denote by $S^{\perp_{\chi}}$ and $S^{\perp_{s}}$ as the sets:
$S^{\perp_{\chi}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{\chi}=0,\,\forall s\in S\}\quad\text{and}\quad S^{\perp_{s}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{s}=0,\,\forall s\in S\}.$
Lemma 3 ([29] Lemma 6).
Let $R$ be a finite commutative chain ring and $S$ be a subgroup of $(R^{2n},+)$. Then $|S|\,|S^{\perp_{\chi}}|=|R^{2n}|$.
Lemma 4 ([29] Lemma 7).
Let $R$ be a finite commutative chain ring and $S$ and $T$ be $R$-submodules of $R^{2n}$. Then $S\perp_{\chi}T$ if and only if $S\perp_{s}T$.
A quantum code $Q$ of length $n$ and size $K$ is a $K$-dimensional subspace of the $\mathfrak{q}^{n}$-dimensional Hilbert space $\mathbb{C}^{\mathfrak{q}^{n}}=\mathbb{C}^{\mathfrak{q}}\otimes\cdots\otimes\mathbb{C}^{\mathfrak{q}}$. A code with a minimum distance $d$ can detect errors that affect up to $d-1$ subsystems non-trivially. Moreover, it can correct errors affecting up to $1/2(d-1)$ of the subsystems. Quantum codes are often denoted with $((n,K,d))_{\mathfrak{q}}$, where $n$, $K$, and $d$ are as defined above. Let $R$ have $\mathfrak{q}$ elements and $B$ be an orthonormal basis of $\mathbb{C}^{\mathfrak{q}}$ concerning the usual Hermitian inner product. The elements of the basis $B$ are labeled by ring elements, such that $B=\{|x\rangle\,|\,x\in R\}$, where ||\,$\rangle$| ⟩ is Dirac ket notation. For each $a\in R$ and $b\in R$, the operator $X(a):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined by $X(a)|x\rangle=|x+a\rangle$, and the operator $Z(b):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined as $Z(b)|x\rangle=\chi(bx)|x\rangle$, for all $x$ in $R$. These operators can be extended for the space $\mathbb{C}^{\mathfrak{q}^{n}}$ as $X(a)=X(a_{1})\otimes X(a_{2})\otimes\cdots\otimes X(a_{n})$ and $Z(b)=Z(b_{1})\otimes Z(b_{2})\otimes\cdots\otimes Z(b_{n})$, where $a=(a_{1},\ldots,a_{n})\in{R}^{n}$ and $b=(b_{1},\ldots,b_{n})\in{R}^{n}$. Let $G_{n}$ be a group consisting of all $\chi(c)X(a)Z(b)$, where $a,b\in R^{n}$ and $c\in R$. For any subgroup $S$ of $G_{n}$, $\text{Fix}(S)$ is defined as
$\text{Fix}(S)=\{v\in\mathbb{C}^{\mathfrak{q}^{n}}\mid Ev=v\text{ for all }E\in S\}.$
The quantum code $Q$ is called stabilizer code if and only if there exists a subgroup $S$ of $G_{n}$ such that $\text{Fix}(S)=Q$.
The symplectic weight of $(a|b)\in R^{2n}$, denoted by $\operatorname{swt}((a|b))$, is defined as the number of indices $i$, where $a_{i}\neq 0$ or $b_{i}\neq 0$. In other words, $\operatorname{swt}((a|b))=\left|\{k\,|\,(a_{k},b_{k})\neq(0,0)\}\right|$. It is well-known that
$\frac{1}{2}\operatorname{wt}((a|b))\leqslant\operatorname{swt}((a|b))\leqslant\operatorname{wt}((a|b)),$ (21)
where $\operatorname{wt}((a|b))$ denotes the Hamming weight of the vector $(a|b)$, see [35].
Theorem 13 ([29] Theorem 8).
An $((n,K,d))_{\mathfrak{q}}$ stabilizer code exists if and only if there exists an additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$ if $K>1$, and $\operatorname{swt}(C^{\perp_{\chi}}\backslash\{0\})=d$ if $K=1$.
A linear code over the ring $R$ of length $n$, rank $k$, and minimum Hamming distance $d$ is denoted by the parameters $[n,k,d]_{R}$. Now, we are ready to present an annihilator counterpart for the Euclidean and Hermitian CSS construction, as presented in [17, 29, 33].
Theorem 14 (Annihilator CSS construction).
Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$. Let $f_{0}\in R$ be an invertible element, $f(x)=g_{1}(x)h_{1}(x)=g_{2}(x)h_{2}(x)$, $C_{1}=\langle g_{1}(x)\rangle$ and $C_{2}=\langle g_{2}(x)\rangle$ be two $f(x)$-polycyclic codes over the ring $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively, such that $C^{\circ}_{2}\subseteq C_{1}$. Then, there exists an $((n,k_{1}+k_{2}-n,d))_{\mathfrak{q}}$ stabilizer code with the minimum distance $d=\min\{\operatorname{wt}(c)\,|\,c\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})\}$.
Proof.
We define $C=C^{\circ}_{1}\times C^{\circ}_{2}\subseteq{R}^{2n}$, where $C_{1}^{\circ}$ and $C_{2}^{\circ}$ are annihilator duals of $C_{1}$ and $C_{2}$, respectively. If $(a|b)$ and $(a^{\prime}|b^{\prime})$ are in $C$, then $b\cdot a^{\prime}-b^{\prime}\cdot a=0$. Thus $C\subseteq C^{\perp_{s}}$, that is $C\perp_{s}C$. Using Lemma 4, we have $C\subseteq C^{\perp_{\chi}}$. Furthermore, if $(a|b)\in C_{2}\times C_{1}$, we have $b\cdot a^{\prime}-b^{\prime}\cdot a=0$ for any arbitrary element $(a^{\prime}|b^{\prime})$ in $C$, and hence $(a|b)\in C^{\perp_{s}}$. So, $(C_{2}\times C_{1})\perp_{s}C$. Applying Lemma 4, $(C_{2}\times C_{1})\perp_{\chi}C$, which implies $(C_{2}\times C_{1})\subseteq C^{\chi}$. On the other side, by Lemma 3 and Part (3) in Theorem 10, we have
$|C^{\perp_{\chi}}|=\frac{|R|^{2n}}{|C|}=\frac{|R|^{n}}{|C^{\circ}_{1}|}\frac{|R|^{n}}{|C^{\circ}_{2}|}=|C_{1}||C_{2}|=|C_{2}\times C_{1}|.$
Therefore, $C_{2}\times C_{1}=C^{\perp_{\chi}}$. Moreover,
$|C|=|C^{\circ}_{1}||C^{\circ}_{2}|=\frac{|R|^{n}}{|C_{1}|}\frac{|R|^{n}}{|C_{2}|}=\frac{|R|^{2n}}{|R|^{k_{1}+k_{2}}}=\frac{|R|^{n}}{|R|^{k_{1}+k_{2}-n}}.$
If we take $K=|{R}|^{k_{1}+k_{2}-n}$, then $|C|=|R|^{n}/K$. We claim that $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$. To prove, assume that $(a|b)\in(C^{\perp_{\chi}}\backslash C)=((C_{2}\times C_{1})\backslash(C^{\circ}_{1}\times C^{\circ}_{2}))$ is an arbitrary element. So, $(a|b)\in C_{2}\times C_{1}$ and $(a|b)\notin C^{\circ}_{1}\times C^{\circ}_{2}$. That is $a\notin C^{\circ}_{1}$ or $b\notin C^{\circ}_{2}$. Therefore, $a,b\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$. According to the definition of $d$, $\operatorname{wt}(a)\geqslant d$ and $\operatorname{wt}(b)\geqslant d$. Applying Equation (21), we have
$\operatorname{swt}(a|b)\geqslant\frac{1}{2}\operatorname{wt}(a|b)=\frac{1}{2}(\operatorname{wt}(a)+\operatorname{wt}(b))\geqslant d.$
Thus, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\geqslant d$. On the other side, we know there exists an element $c$ in $(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$ such that $\operatorname{wt}(c)=d$. In the first case, let $c\in(C_{1}\backslash C_{2}^{\circ})$. Then we have $(0|c)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(0|c)=\operatorname{wt}(c)=d$. In the second case, let $c\in(C_{2}\backslash C_{1}^{\circ})$. Then we have $(c|0)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(c|0)=\operatorname{wt}(c)=d$. Hence in both cases, there exists an element in $(C^{\perp_{\chi}}\backslash C)$ with the symplectic weight $d$. Therefore, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\leqslant d$. To sum up, we have found the additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{wt}(C^{\perp_{\chi}}\backslash C)=d$ As a result, Theorem 13 implies that there exists an $((n,K,d))_{\mathfrak{q}}$ stabilizer code. ∎
Corollary 4.
Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$. Let $f_{0}\in R$ be an invertible element and $f(x)=g(x)h(x)$. If $C$ is an $f(x)$-polycyclic code over the chain ring $R$ with parameters $[n,k,d]_{R}$, generated by $g(x)$, and $C^{\circ}\subseteq C$, then there exists an $((n,2k-n,\geqslant d))_{\mathfrak{q}}$ stabilizer code.
Now, we are prepared to give the result that allows to construct quantum codes from $f(x)$-polycyclic codes over the serial ring $\mathscr{R}$.
Theorem 15.
Suppose that $f(x)$ defined as in Equation (3) is a polynomial in $\mathscr{R}[x]$, $f_{0}\in\mathscr{R}$ is an invertible element, and $f(x)=g(x)h(x)$. Let $\mathscr{C}$ be an annihilator dual-preserving $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$, decomposed as $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}$. If each $\mathscr{C}_{i}$ has parameters $[n_{i},k_{i},d_{i}]$, then there exists a stabilizer code with parameters $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$.
Proof.
By Corollary 3, $\mathscr{C}_{i}$ is dual-preserving $f_{i}(x)$-polycyclic code over the chain ring $\mathscr{R}_{i}$ for all $i$. Suppose that $\big{(}g_{1}(x),g_{2}(x),\ldots,g_{k}(x)\big{)}$ and $\big{(}h_{1}(x),h_{2}(x)\ldots,h_{k}(x)\big{)}$ are the images of $g(x)$ and $h(x)$ under the mapping $\eta$ defined in Eq. (7), respectively. Since $g(x)=\oplus_{i=1}^{k}g_{i}(x)e_{i}$ and $h(x)=\oplus_{i=1}^{k}h_{i}(x)e_{i}$, we have $f_{i}(x)=g_{i}(x)h_{i}(x)$ for all $i$. Moreover, since $f_{0}$ is invertible, the constant term of the polynomial $f_{i}(x)$ is invertible for all $i$. By Corollary 4, there exists an $((n_{i},2k_{i}-n_{i},\geqslant d_{i}))$ stabilizer code for all $i$. It is well-know that the direct sum of two quantum-error-correcting codes $((n,k,d))_{\mathfrak{q}}$ and $((n^{\prime},k^{\prime},d^{\prime}))_{\mathfrak{q}}$ is an $((n+n^{\prime},k+k^{\prime},d^{\prime\prime}))_{\mathfrak{q}}$ code, where $d^{\prime\prime}=\min\{d,d^{\prime}\}$, see [8] Section 4. Therefore, there exists a $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$ stabilizer code, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$. ∎
8 Conclusion
In this paper, we have investigated the structure of $f(x)$-polycyclic codes over a specific category of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial. We have established a unified framework in the literature of constacyclic and polycyclic codes over various rings, all of which are examples of the ring $\mathscr{R}$. We have defined the new concepts QsDP codes and $f(x)$-polycyclic-QsDP codes, and have proved that $f(x)$-polycyclic codes over $\mathscr{R}$ are isomorphic to $f(x)$-polycyclic-QsDP codes. Additionally, we have established a polynomial structure for the Eclidean dual of $f(x)$-polycyclic codes over $\mathscr{R}$. The annihilator dual is an alternative duality concept ensuring that the dual of $f(x)$-polycyclic codes has an $f(x)$- polycyclic structure, a feature that fails in Euclidean dual. We have studied some properties of the annihilator dual of $f(x)$-polycyclic code over the ring $\mathscr{R}$ and have provided necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator LCD, and annihilator dual-containing polycyclic codes over $\mathscr{R}$. Finally, we have provided an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.

### v3

@@
 We provide necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator linear complementary dual, and annihilator dual-containing polycyclic codes over this class of rings.
-We also establish the CSS construction for annihilator dual-preserving polycyclic codes over the chain ring $R$ and use this construction to derive quantum codes from polycyclic codes over $\mathscr{R}$.
+We also establish the CSS construction for annihilator dual-preserving polycyclic codes over Frobenius rings.
 keywords:
 Polycyclic code, Serial ring, Annihilator dual, quasi-$s$-dimensional polycyclic code, annihilator CSS construction AMS Subject Classification 2010: 94B15, 13B25, 81P70
-journal: Cryptography and Communicationsmmfootnotetext: Maryam Bajalan is supported by the Bulgarian Ministry of Education and Science, Scientific Programme ”Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No.
-DO1-67/05.05.2022.vvfootnotetext: Both authors were partially supported by Grant TED2021-130358B-I00 funded by MCIU/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”
+journal: Cryptography and Communicationsmmfootnotetext: Maryam Bajalan is supported by the Bulgarian Ministry of Education and Science, Scientific Programme “Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No.
+DO1-67/05.05.2022.vvfootnotetext: Both authors were partially supported by Grant TED2021-130358B-I00 funded by MCIU/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”.
+This research was partially conducted during a research stay of the second author funded by the Institute of Mathematics and Informatics at the Bulgarian Academy of Sciences.
 1 Introduction
 Polycyclic codes are ideals in the quotient of the polynomial ring $A[x]/\langle f(x)\rangle$, where $A$ is a ring and $f(x)$ is a polynomial in $A[x]$, see [1, 8, 9, 28].
-Cyclic codes and constacyclic codes over $A$ are particular instances when $f(x)=x^{n}-1$ and $f(x)=x-\lambda$, respectively.
+Cyclic codes and constacyclic codes over $A$ are particular instances, when $f(x)=x^{n}-1$ and $f(x)=x-\lambda$, respectively.
 A ring is called uniserial if all its submodules are linearly ordered by inclusion, and it is called serial if it is a direct sum of uniserial rings.
 In this paper, we aim to investigate polycyclic codes in the case $A$ is the serial ring $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial for $i=1,2,\ldots,s$.
-There is a vast literature of polycyclic codes over some special classes of rings, that we slightly review here.
+We briefly review the vast literature on polycyclic codes over some special classes of rings.
 Alkenani et al.
-[3] studyed constacyclic codes over the ring $\mathbb{F}_{q}[u_{1},u_{2}]/\langle u_{1}^{2}-u_{1},u_{2}^{2}-u_{2},u_{1}u_{2}-u_{2}u_{1}\rangle$.
+[3] studied constacyclic codes over the ring $\mathbb{F}_{q}[u_{1},u_{2}]/\langle u_{1}^{2}-u_{1},u_{2}^{2}-u_{2},u_{1}u_{2}-u_{2}u_{1}\rangle$.
 Dinh et al.
@@
 Qi [37] focused on polycyclic codes over $\mathbb{F}_{q}+u\mathbb{F}_{q}$, with $u^{2}=0$.
-Karthick [22] enquired about polycyclic codes over the ring $R=\mathbb{F}_{q}+u\mathbb{F}_{q}+v\mathbb{F}_{q}$, subject to the conditions $u^{2}=\alpha u$, $v^{2}=v$, and $uv=vu=0$, where $\alpha$ represents a unit element of $R$.
+Karthick [26] investigated polycyclic codes over the ring $R=\mathbb{F}_{q}+u\mathbb{F}_{q}+v\mathbb{F}_{q}$, subject to the conditions $u^{2}=\alpha u$, $v^{2}=v$, and $uv=vu=0$, where $\alpha$ represents a unit element of $R$.
 The serial ring $\mathscr{R}=\frac{R[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}$, where $R$ is a finite commutative chain ring with maximal ideal $\mathfrak{m}$ and the natural projection of $t_{i}$ on $(R/\mathfrak{m})[x_{i}]$ is a square free polynomial for each $i\in\{1,\ldots,s\}$, is a comprehensive generalization of the rings previously discussed in the papers [2, 3, 6, 10, 14, 19, 20, 25, 26, 37].
@@
 In Section 5, using multivariable codes over chain rings in [29, 30], we establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case where $f(x)$ is a polynomial with coefficients in $R$.
-The annihilator dual is an alternative duality concept ensuring that the dual of polycyclic codes has a polycyclic structure, a feature that fails in Euclidean dual.
+The annihilator dual is an alternative duality concept ensuring that the dual of polycyclic codes has a polycyclic structure, a feature that fails in the Euclidean dual.
 In Section 6, some properties of the annihilator dual of a polycyclic code over $\mathscr{R}$ are studied.
@@
 Existing literature has applied this construction only to codes that preserve Euclidean or Hermitian duality.
-In this paper, we present construction based on preserving the annihilator duality.
-In Section 7, we provide an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.
+In Section 7, we present this construction for polycyclic codes over Frobenius rings that preserve annihilator duality.
+List of symbols and abbreviations
+  • $R$
+    : Finite chain ring with the maximal ideal $\mathfrak{m}=\langle\gamma\rangle$ and the nilpotency index $e$.
+  • $\mathbb{F}_{q}$
+    : Residue field $\mathbb{F}_{q}=R/\mathfrak{m}$, where $q=p^{t}$, $p$ a prime number.
+  • $\overline{\phantom{r}}$
+    : Natural ring homomorphism from $R[y]$ to $\mathbb{F}_{q}[y]$.
+  • $\mathscr{R}$
+    : Ring $R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $t_{i}(x_{i})\in R[x_{i}]$ is a monic square-free of degree $m_{i}$.
+  • $H_{i}$
+    : Set of roots of $\bar{t}_{i}(x_{i})$ in an extension of $\mathbb{F}_{q}$.
+  • $C(\nu)$
+    : Set of the class of $\nu\in\mathcal{H}=\prod_{i=1}^{s}H_{i}$ as $\{(\nu^{q^{j}_{1}},\ldots,\nu^{q^{j}_{s}})\mid j\in\mathbb{N}\}$.
+  • $\mathcal{C}$
+    : Set of all classes $C(\nu)$.
+  • $C$
+    : Simpler notation for $C(\nu)$.
+  • $p_{C,i}(x_{i})$
+    : Polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$.
+  • $b_{C,i}(x_{i})$
+    : Polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))$.
+  • $\widetilde{b}_{C,i}(x_{i})$
+    : Polynomial $p_{C,i}(x_{i})/b_{C,i}(x_{i})$.
+  • $w_{C,i}(x_{1},\ldots,x_{i})$
+    : Polynomial obtained by replacing $\nu_{i}$ with $x_{i}$ in $b_{C,i}(x_{i})$.
+  • $\pi_{C,i}(x_{1},\ldots,x_{i})$
+    : Polynomial obtained by replacing $\nu_{i}$ by $x_{i}$ in $\widetilde{b}_{C,i}(x_{i})$.
+  • $q_{C,i}$
+    : Hensel lift to $R$ of $p_{C,i}$.
+  • $z_{C,i}$
+    : Hensel lift to $R$ of $w_{C,i}$.
+  • $\sigma_{C,i}$
+    : Hensel lift to $R$ of $\pi_{C,i}$.
+  • $I_{C}$
+    : Ideal $\langle q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle$.
+  • $T_{C}$
+    : Chain ring $R[x_{1},\ldots,x_{s}]/I_{C}$.
+  • $h_{C}$
+    : Polynomial $h_{C}(x_{1},x_{2},\ldots,x_{s})=\prod_{i=1}^{s}t_{i}(x_{i})q_{C,i}(x_{i})\prod_{i=1}^{s}\sigma_{q_{C,i}}(x_{2},\ldots,x_{i})$.
+  • $e_{C}$
+    : Orthogonal idempotent such that $\mathscr{R}e_{C}\cong\langle h_{C}+I\rangle$.
+  • $\mathscr{R}_{i}$
+    : Chain ring isomorphic to $\mathscr{R}e_{C_{i}}$, where $\{C_{i}\}_{i=1}^{k}$ is a set of representatives of $\mathcal{C}$.
+  • $\mathscr{C}$
+    : Linear code of length $n$ over $\mathscr{R}$.
+  • $f(x)$
+    : Polynomial $x^{n}-(f_{n-1}x^{n-1}+\cdots+f_{1}x+f_{0})$ in $\mathscr{R}[x]$.
+  • $\tau_{f(x)}$
+    : $f(x)$-polycyclic shift.
+  • $\boldsymbol{c}$
+    : Element $(c_{0},c_{1},\ldots,c_{n-1})\in\mathscr{R}^{n}$.
+  • $c_{i,j}$
+    : Components of $c_{j}$ ($10\leqslant j\leqslant n-$ 1) in $c_{j}=c_{1,j}e_{C_{1}}+c_{2,j}e_{C_{2}}+\cdots+c_{k,j}e_{C_{k}}\in\mathscr{R}$.
+  • $\boldsymbol{c}_{i}$
+    : Components of $\boldsymbol{c}$ in $\boldsymbol{c}=\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}$.
+  • $\mathscr{C}_{i}$
+    : Set $\{\boldsymbol{c}_{i}\in\mathscr{R}^{n}\mid\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}\in\mathscr{C}\}$.
+  • $f_{i}(x)$
+    : Polynomial $x^{n}-(f_{i,n-1}x^{n-1}+\cdots+f_{i,1}x+f_{i,0})$, where $f_{i,j}$ is component of $f_{j}$.
+  • $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$
+    : Minimum number of generators for spanning $\mathscr{C}$ as $\mathscr{R}$-linear combinations.
+  • $\mathscr{R}_{f}$
+    : Quotient ring $\mathscr{R}[x]/\langle f(x)\rangle$.
+  • $\phi$
+    : Map that associates each vector in $\mathscr{R}$ with its corresponding polynomial in $\mathscr{R}_{f}$.
+  • $\mathbf{x}^{\alpha}$
+    : Monomial $x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\ldots x_{s}^{\alpha_{s}}$, where $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{s})\in(\mathbb{N}\cup\{0\})^{s}$.
+  • $\succ$
+    : Lexicographic ordering on the monomials.
+  • $\Phi$
+    : $R$-isomorphism from $\mathscr{R}$ to $R^{m}$, where $m=m_{1}m_{2}\ldots m_{s}$.
+  • $QsDP$
+    : Quasi-s-dimensional polycyclic codes.
+  • $C_{f}$
+    : Companion matrix of the polynomial $f$.
+  • $I_{i\times i}$
+    : Identity matrix.
+  • $\otimes$
+    : Tensor product of matrices.
+  • $\nu_{f(x)}$
+    : $f(x)$-sequential shift.
+  • $\mathscr{C}^{\perp}$
+    : Euclidean dual of $\mathscr{C}$.
+  • $\mathscr{C}^{\circ}$
+    : Annihilator dual of $\mathscr{C}$.
+  • $LCD$
+    : Linear complementary dual code.
+  • $\chi$
+    : Character in $\operatorname{Hom}(R;\mathbb{C}^{*})$.
+  • $\operatorname{Ann}(\mathscr{C})$
+    : Annihilator of the subset $\mathscr{C}$ of $\mathscr{R}_{f}$.
+  • $\operatorname{swt}$
+    : Symplectic weight.
+  • $\operatorname{wt}$
+    : Hamming weight.
 2 The structure of the ring $\mathscr{R}=\frac{R[x_{1},\ldots,x_{s}]}{\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle}$
@@
 Additionally, it is called regular if it is not a zero divisor in $R[y]$.
-In [11, Proposition 2.3] it is stated that $p(y)$ is regular if and only if $p_{i}$ is a unit element for some $i$.
+In [15, Proposition 2.3] it is stated that $p(y)$ is regular if and only if $p_{i}$ is a unit element for some $i$.
 Hence, monic polynomials are regular.
-If $g(y)\in R[y]$ is a regular basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ forms a chain ring, see [11, Lemma 3.1].
-It is well-known that the monic square-free polynomial $g(y)$ factors uniquely as a product of monic basic irreducible pairwise coprime polynomials, given by $g(y)=g_{1}(y)g_{2}(y)\cdots g_{r}(y)$[28, Theorem XIII.4].
+If $g(y)\in R[y]$ is a regular basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ forms a chain ring, see [15, Lemma 3.1].
+The monic square-free polynomial $g(y)$ factors uniquely as a product of monic basic irreducible pairwise coprime polynomials, given by $g(y)=g_{1}(y)g_{2}(y)\cdots g_{r}(y)$ [33, Theorem XIII.4].
 Hence, by the Chinese Remainder Theorem, the quotient ring $R[y]/\langle g(y)\rangle$ can be decomposed as a direct sum of chain rings
@@
 Let $p_{C,i}(x_{i})$ denote the polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$ and $C=C(\nu)$ is a class of $\nu$.
-For each $i=2,\ldots,s$, let $b_{C,i}(x_{i})=\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))\in\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1})[x_{i}]$ and $\widetilde{b}_{C,i}(x_{i})=\frac{p_{C,i}(x_{i})}{b_{C,i}(x_{i})}$.
+For each $i=2,\ldots,s$, let $b_{C,i}(x_{i})=\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))\in\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1})[x_{i}]$ and $\widetilde{b}_{C,i}(x_{i})=p_{C,i}(x_{i})/b_{C,i}(x_{i})$.
 The polynomials $b_{C,i}(x_{i})$ and $\widetilde{b}_{C,i}(x_{i})$ are coprime polynomials.
@@
   4. 4.
-    If $\lvert\mathcal{C}\rvert=k$ and $\{C_{i}\}_{i=1}^{k}$ is a set of representatives $\mathcal{C}$, then the ring $\mathscr{R}$ can be decomposed as a direct sum
+    If $\lvert\mathcal{C}\rvert=k$ and $\{C_{i}\}_{i=1}^{k}$ is a set of representatives of $\mathcal{C}$, then the ring $\mathscr{R}$ can be decomposed as a direct sum
     $\mathscr{R}\cong\bigoplus_{i=1}^{k}\mathscr{R}e_{C_{i}}\cong\bigoplus_{i=1}^{k}\mathscr{R}_{i},$
@@
 A ring is called serial if it is a direct sum of uniserial rings.
-Any uniserial ring is local, and a serial ring is uniserial if and only if it is local, see [31, Section 1.4].
+Any uniserial ring is local, and a serial ring is uniserial if and only if it is local, see [36, Section 1.4].
 Note that any chain ring is a serial ring, thus, if $g(y)\in R[y]$ is a monic basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ is a serial ring.
@@
 Now, we define the ring isomorphism
-    $\eta:\mathscr{R}_{f}=\frac{\mathscr{R}[x]}{\langle f(x)\rangle}$ $\longrightarrow\frac{\mathscr{R}_{1}[x]}{\langle f_{1}(x)\rangle}\times\frac{\mathscr{R}_{2}[x]}{\langle f_{2}(x)\rangle}\times\dots\times\frac{\mathscr{R}_{k}[x]}{\langle f_{k}(x)\rangle},$ (7)
+    $\eta:\mathscr{R}_{f}=\frac{\mathscr{R}[x]}{\langle f(x)\rangle}$ $\longrightarrow\frac{\mathscr{R}_{1}[x]}{\langle f_{1}(x)\rangle}\oplus\frac{\mathscr{R}_{2}[x]}{\langle f_{2}(x)\rangle}\oplus\dots\oplus\frac{\mathscr{R}_{k}[x]}{\langle f_{k}(x)\rangle},$ (7)
     $g(x)$ $\longmapsto\big{(}g_{1}(x),\,g_{2}(x),\,\dots,\,g_{k}(x)\big{)}.$
@@
   • 1.
-    The condition for the ring $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ being a principal ideal ring in terms of the chain ring $\mathscr{R}_{i}$ and the polynomial $f_{i}(x)$ can be found in [26, Theorem 1].
+    The condition for the ring $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ being a principal ideal ring in terms of the chain ring $\mathscr{R}_{i}$ and the polynomial $f_{i}(x)$ can be found in [31, Theorem 1].
   • 2.
-    If $f_{i}(x)$ is a regular basic irreducible polynomial in the ring $\mathscr{R}_{i}[x]$, then $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma_{i}+\langle f_{i}(x)\rangle\rangle$, where $\gamma_{i}$ denotes the maximal ideal of the chain ring $\mathscr{R}_{i}$, see [11, Lemma 3.1].
+    If $f_{i}(x)$ is a regular basic irreducible polynomial in the ring $\mathscr{R}_{i}[x]$, then $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma_{i}+\langle f_{i}(x)\rangle\rangle$, where $\gamma_{i}$ denotes the maximal ideal of the chain ring $\mathscr{R}_{i}$, see [15, Lemma 3.1].
 Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$, we will denote the rank of the code $\mathscr{C}$ as $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$, i.e.
-the minimum number of generators needed to span $\mathscr{C}$ as $\mathscr{R}$-linear combinations.
+the minimum number of generators for spanning $\mathscr{C}$ as $\mathscr{R}$-linear combinations.
 We will also denote the minimum Hamming distance of $\mathscr{C}$ as $d(\mathscr{C})$.
@@
     $d(\mathscr{C})=\min\{d(\mathscr{C}_{i})\mid i=1,\ldots,k\}$.
-  8. 8.
-    $\mathscr{C}$ is a free over $\mathscr{R}$ if and only if every $\mathscr{C}_{i}$ is free over $\mathscr{R}_{i}$ for $i=\{1,\ldots,k\}$.
 Proof.
@@
 Therefore, $d_{\star}=d(\mathscr{C})$.
-  8. 8.
-[…]
 4 Description of polycyclic codes over $\mathscr{R}$ as $R$-submodules
@@
 Therefore, $\Psi(c)$ consist of $n$-subarrays arranged as follows:
-($\begin{matrix}c^{0,0}_{0}&\cdots&c^{m_{1}-1,0}_{0}\\ c^{0,1}_{0}&\cdots&c^{m_{1}-1,1}_{0}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{0}&\cdots&c^{m_{1}-1,m_{2}-1}_{0}\\ \end{matrix}&\begin{matrix}c^{0,0}_{1}&\cdots&c^{m_{1}-1,0}_{1}\\ c^{0,1}_{1}&\cdots&c^{m_{1}-1,1}_{1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{1}&\cdots&c^{m_{1}-1,m_{2}-1}_{1}\\ \end{matrix}&\,\,\cdots&\begin{matrix}c^{0,0}_{n-1}&\cdots&c^{m_{1}-1,0}_{n-1}\\ c^{0,1}_{n-1}&\cdots&c^{m_{1}-1,1}_{n-1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{n-1}&\cdots&c^{m_{1}-1,m_{2}-1}_{n-1}\\ \end{matrix}\\ \end{array}\end{pmatrix}.$ (12)
+$\begin{pmatrix}\begin{array}[]{c|c|c|c}\begin{matrix}c^{0,0}_{0}&\cdots&c^{m_{1}-1,0}_{0}\\ c^{0,1}_{0}&\cdots&c^{m_{1}-1,1}_{0}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{0}&\cdots&c^{m_{1}-1,m_{2}-1}_{0}\\ \end{matrix}&\begin{matrix}c^{0,0}_{1}&\cdots&c^{m_{1}-1,0}_{1}\\ c^{0,1}_{1}&\cdots&c^{m_{1}-1,1}_{1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{1}&\cdots&c^{m_{1}-1,m_{2}-1}_{1}\\ \end{matrix}&\,\,\cdots&\begin{matrix}c^{0,0}_{n-1}&\cdots&c^{m_{1}-1,0}_{n-1}\\ c^{0,1}_{n-1}&\cdots&c^{m_{1}-1,1}_{n-1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{n-1}&\cdots&c^{m_{1}-1,m_{2}-1}_{n-1}\\ \end{matrix}\\ \end{array}\end{pmatrix}.$ (12)
 Remark 2.
@@
 Let $m=m_{1}m_{2}$.
-A length $mn$ linear code $C$ over the chain ring $R$ with codewords given as in Equation (12) is called quasi-2-dimensional polycyclic (Q2DP) code if the columns of each subarray are closed under $t_{1}(x_{1})$-polycyclic shift and the rows of each subarray are closed under $t_{2}(x_{2})$-polycyclic shift.
+A length $mn$ linear code $C$ over the chain ring $R$ with codewords given as in Equation (12) is a called quasi-2-dimensional polycyclic (Q2DP) code if the columns of each subarray are closed under $t_{1}(x_{1})$-polycyclic shift and the rows of each subarray are closed under $t_{2}(x_{2})$-polycyclic shift.
 The above definition is a generalization of quasi-2-dimensional cyclic (Q2DC) codes defined in [22]. Similarly, we can define a multidimensional version of Q2DP codes and abbreviate them as QsDP.
@@
 Similarly, the above theorem and corollary are true for any value of $s$.
-It is worth noting that $\{1,x_{1},x_{2},x_{1}x_{2}\}$ forms the basis for $\mathscr{R}$, and the representation matrices for these basis elements are $\{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0},C_{t_{2}}^{0}\otimes C_{t_{1}},C_{t_{2}}\otimes C_{t_{1}}^{0},C_{t_{2}}\otimes C_{t_{1}}\}$, see [5, Section 6].
+It is worth noting that $\{1,x_{1},x_{2},x_{1}x_{2}\}$ forms the basis for $\mathscr{R}$, and the representation matrices for these basis elements are $\{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0},C_{t_{2}}^{0}\otimes C_{t_{1}},C_{t_{2}}\otimes C_{t_{1}}^{0},C_{t_{2}}\otimes C_{t_{1}}\}$, see [8, Section 6].
 We notice that the matrix $M$ is constructed based on this set.
-Therefore, for the general case $s$, the matrix $M$ can be constructed using the representation matrices of the basis elements of $\mathscr{R}$ and $R$-module.
+Therefore, for the general case $s$, the matrix $M$ can be constructed using the representation matrices of the basis elements of $\mathscr{R}$ and $R$-modules.
 5 Euclidean dual of polycyclic codes over $\mathscr{R}$
@@
 It is easy to see that the operator $\nu_{f(x)}$ is the adjoint operator of $\tau_{f(x)}$, that is $\langle\tau_{f(x)}(x),y\rangle=\langle x,\nu_{f(x)}(y)\rangle$.
-It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator Therefore, we have the following theorem.
+It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator.
+Therefore, we have the following theorem.
 Theorem 5.
@@
 Theorem 6.
-[23, Theorem 3.5] The codes $\mathscr{C}$ and $\mathscr{C}^{\perp}$ are $f(x)$-polycyclic codes if and only if they are both $f(x)$-sequential codes, if and only if $\mathscr{C}$ is a constacyclic code.
+[28, Theorem 3.5] The codes $\mathscr{C}$ and $\mathscr{C}^{\perp}$ are $f(x)$-polycyclic codes if and only if they are both $f(x)$-sequential codes, if and only if $\mathscr{C}$ is a constacyclic code.
 Using the results in [29], we can establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case of $f(x)$.
@@
 The set of vectors $\{e_{i}:1\leqslant i\leqslant n\}$, where each $e_{i}$ has a value of $1$ in the $i$-th position and 00 elsewhere, forms a basis for $\mathscr{R}^{n}$.
-The Gram matrix $A$ associated with this set is a matrix with entries $a_{i,j}=\langle e_{i},e_{j}\rangle_{f}$.
-For a linear code $\mathscr{C}$ over $\mathscr{R}$, we define the set $\mathscr{C}\cdot A:=\{cA\mid c\in\mathscr{C}\}$.
+The Gram matrix $A$ associated with this set and the polynomial $f$ is a matrix with entries $a_{i,j}=\langle e_{i},e_{j}\rangle_{f}$.
+For a linear code $\mathscr{C}$ over $\mathscr{R}$, we define the set $\mathscr{C}A:=\{cA\mid c\in\mathscr{C}\}$.
 Similar to Lemma 4 and Corollary 1 in [18], we have the following lemma.
@@
   1. 1.
-    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$.
+    $\mathscr{C}^{\circ}=(\mathscr{C}A)^{\perp}$.
   2. 2.
@@
     First, we claim that the annihilator product defined on each $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is non-degenerate.
-To prove it, note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the finite ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$, see [28, Chapter 1].
+To prove it, note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the finite ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$, see [33, Chapter 1].
 Thus, since the constant part of the polynomial $f(x)$ is invertible, the constant part of each polynomial $f_{i}(x)$ is also invertible, and therefore, by Theorem 10 the claim follows.
@@
 Consequently, by Lemma 2, we have
-    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$ $=\bigg{(}\big{(}\bigoplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\big{)}\cdot\big{(}\bigoplus_{i=1}^{k}A_{i}e_{C_{i}}\big{)}\bigg{)}^{\perp}$
-    $=\bigg{(}\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}e_{C_{i}}\bigg{)}^{\perp}$
-    $=\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}^{\perp}e_{C_{i}}=\bigoplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}.$
+    $\mathscr{C}^{\circ}=\left(\mathscr{C}A\right)^{\perp}$ $=\left(\bigg{(}\bigoplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\bigg{)}\bigg{(}\bigoplus_{i=1}^{k}A_{i}e_{C_{i}}\bigg{)}\right)^{\perp}$
+    $=\left(\bigoplus_{i=1}^{k}\bigg{(}\mathscr{C}_{i}A_{i}\bigg{)}e_{C_{i}}\right)^{\perp}$
+    $=\bigoplus_{i=1}^{k}\bigg{(}\mathscr{C}_{i}A_{i}\bigg{)}^{\perp}e_{C_{i}}=\bigoplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}.$
   2. 2.
@@
 Theorem 12.
-Suppose Let $f_{0}\in\mathscr{R}$ be an invertible element, $\mathscr{C}$ is an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ generated by a single polynomial $g(x)\in\mathscr{R}_{f}$, and the leading coefficient of $g(x)$ be a unit element in $\mathscr{R}$ and there exist $h(x)\in\mathscr{R}_{f}$ such that $f(x)=g(x)h(x)$.
+Let $f_{0}\in\mathscr{R}$ be an invertible element, $\mathscr{C}$ is an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ generated by a single polynomial $g(x)\in\mathscr{R}_{f}$, and the leading coefficient of $g(x)$ be a unit element in $\mathscr{R}$ and there exist $h(x)\in\mathscr{R}_{f}$ such that $f(x)=g(x)h(x)$.
 Then
@@
 We will proof the first statement, and the other ones follow similarly.
-Note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$ [28, Chapter 1].
+Note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$ [33, Chapter 1].
 Thus, if the leading coefficient of $g(x)$ is a unit element in $\mathscr{R}$, then the leading coefficient of $g_{i}(x)$ is a unit element in $\mathscr{R}_{i}$ for all $i\in\{1,\ldots,k\}$.
@@
 Consider the ring $\mathscr{R}=R[x_{1},x_{2},x_{3}]/\langle x_{1}^{2}-1,x_{2}^{2}-1,x_{3}^{2}-1\rangle$, where $R$ is a chain ring with the residue field $\mathbb{F}_{q}$.
-Any element $r\in\mathscr{R}$ of the form $r=a_{1}+a_{2}x_{1}+a_{3}x_{2}+a_{4}x_{3}+a_{5}x_{1}x_{2}+a_{6}x_{2}x_{3}+a_{7}x_{1}x_{3}+a_{8}x_{1}x_{2}x_{3}$ can be uniquely expressed as $r_{1}e_{1}+r_{2}e_{2}+r_{3}e_{3}+r_{4}e_{4}+r_{5}e_{5}+r_{6}e_{6}+r_{7}e_{7}+e_{8}e_{8}$, where $r_{i}\in\mathscr{R}$ and
+With a similar approach as in [4], any element $r\in\mathscr{R}$ of the form $r=a_{1}+a_{2}x_{1}+a_{3}x_{2}+a_{4}x_{3}+a_{5}x_{1}x_{2}+a_{6}x_{2}x_{3}+a_{7}x_{1}x_{3}+a_{8}x_{1}x_{2}x_{3}$ can be uniquely expressed as $r_{1}e_{1}+r_{2}e_{2}+r_{3}e_{3}+r_{4}e_{4}+r_{5}e_{5}+r_{6}e_{6}+r_{7}e_{7}+r_{7}e_{8}$, where $r_{i}\in\mathscr{R}$ and
 $e_{1}$ $=t(1+x_{1}+x_{2}+x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}+x_{1}x_{2}x_{3}),$
@@
 $e_{8}$ $=t(1-x_{1}-x_{2}-x_{3}+x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}-x_{1}x_{2}x_{3}),$
-where $t\in\mathbb{F}_{q}$ and $8t=1\mod p$. As special case, let $R=\mathbb{Z}_{9}$ and $t=2$.
+where $t\in\mathbb{F}_{q}$ and $8t=1\mod p$. As a special case, let $R=\mathbb{Z}_{9}$ and $t=2$.
   1. 1.
@@
 7 Annihilator CSS construction
-The CSS construction is a well-known method to construct quantum codes from linear codes over fields or rings, see for example [17, 29, 33].
-Depending on the duality defined for the code, this technique is referred to as either the Euclidean CSS construction or the Hermitian CSS construction.
-As we noticed above, polycyclic codes are also equipped with the annihilator dual.
-In this section, we establish annihilator CSS construction, a novel approach for constructing quantum codes from annihilator dual-preserving polycyclic codes over finite chain rings.
-Recall that since $R$ is a finite commutative chain ring with identity it is a Frobenius ring, thus it has generating characters.
-For the character $\chi$ in $\operatorname{Hom}(R;\mathbb{C}^{*})$, there exists a unique function $\psi$ in $\operatorname{Hom}(R;\mathbb{Q}/\mathbb{Z})$ such that $\chi(x)=\exp(2\pi i\psi(x))$.
-Let $\langle\,,\,\rangle_{\chi}:{R}^{2n}\times{R}^{2n}\rightarrow\mathbb{Q}/\mathbb{Z}$ and $\langle\,,\,\rangle_{s}:{R}^{2n}\times{R}^{2n}\rightarrow{R}$ be two forms defined as follows:
-$\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{\chi}=\psi(b\cdot a^{\prime}-b^{\prime}\cdot a)\quad\text{and}\quad\langle(a|b),(a^{\prime}|b^{\prime})\rangle_{s}=b\cdot a^{\prime}-b^{\prime}\cdot a.$
-We denote by $S^{\perp_{\chi}}$ and $S^{\perp_{s}}$ as the sets:
-$S^{\perp_{\chi}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{\chi}=0,\,\forall s\in S\}\quad\text{and}\quad S^{\perp_{s}}=\{u\in{R}^{2n}\,|\,\langle(s,u)\rangle_{s}=0,\,\forall s\in S\}.$
-Lemma 3 ([29] Lemma 6).
-Let $R$ be a finite commutative chain ring and $S$ be a subgroup of $(R^{2n},+)$. Then $|S|\,|S^{\perp_{\chi}}|=|R^{2n}|$.
-Lemma 4 ([29] Lemma 7).
-Let $R$ be a finite commutative chain ring and $S$ and $T$ be $R$-submodules of $R^{2n}$. Then $S\perp_{\chi}T$ if and only if $S\perp_{s}T$.
-A quantum code $Q$ of length $n$ and size $K$ is a $K$-dimensional subspace of the $\mathfrak{q}^{n}$-dimensional Hilbert space $\mathbb{C}^{\mathfrak{q}^{n}}=\mathbb{C}^{\mathfrak{q}}\otimes\cdots\otimes\mathbb{C}^{\mathfrak{q}}$.
-A code with a minimum distance $d$ can detect errors that affect up to $d-1$ subsystems non-trivially.
-Moreover, it can correct errors affecting up to $1/2(d-1)$ of the subsystems.
-Quantum codes are often denoted with $((n,K,d))_{\mathfrak{q}}$, where $n$, $K$, and $d$ are as defined above.
-Let $R$ have $\mathfrak{q}$ elements and $B$ be an orthonormal basis of $\mathbb{C}^{\mathfrak{q}}$ concerning the usual Hermitian inner product.
-The elements of the basis $B$ are labeled by ring elements, such that $B=\{|x\rangle\,|\,x\in R\}$, where ||\,$\rangle$| ⟩ is Dirac ket notation.
-For each $a\in R$ and $b\in R$, the operator $X(a):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined by $X(a)|x\rangle=|x+a\rangle$, and the operator $Z(b):\mathbb{C}^{\mathfrak{q}}\rightarrow\mathbb{C}^{\mathfrak{q}}$ is defined as $Z(b)|x\rangle=\chi(bx)|x\rangle$, for all $x$ in $R$.
-These operators can be extended for the space $\mathbb{C}^{\mathfrak{q}^{n}}$ as $X(a)=X(a_{1})\otimes X(a_{2})\otimes\cdots\otimes X(a_{n})$ and $Z(b)=Z(b_{1})\otimes Z(b_{2})\otimes\cdots\otimes Z(b_{n})$, where $a=(a_{1},\ldots,a_{n})\in{R}^{n}$ and $b=(b_{1},\ldots,b_{n})\in{R}^{n}$.
-Let $G_{n}$ be a group consisting of all $\chi(c)X(a)Z(b)$, where $a,b\in R^{n}$ and $c\in R$.
-For any subgroup $S$ of $G_{n}$, $\text{Fix}(S)$ is defined as
-$\text{Fix}(S)=\{v\in\mathbb{C}^{\mathfrak{q}^{n}}\mid Ev=v\text{ for all }E\in S\}.$
-The quantum code $Q$ is called stabilizer code if and only if there exists a subgroup $S$ of $G_{n}$ such that $\text{Fix}(S)=Q$.
-The symplectic weight of $(a|b)\in R^{2n}$, denoted by $\operatorname{swt}((a|b))$, is defined as the number of indices $i$, where $a_{i}\neq 0$ or $b_{i}\neq 0$.
-In other words, $\operatorname{swt}((a|b))=\left|\{k\,|\,(a_{k},b_{k})\neq(0,0)\}\right|$.
-It is well-known that
-$\frac{1}{2}\operatorname{wt}((a|b))\leqslant\operatorname{swt}((a|b))\leqslant\operatorname{wt}((a|b)),$ (21)
-where $\operatorname{wt}((a|b))$ denotes the Hamming weight of the vector $(a|b)$, see [35].
-Theorem 13 ([29] Theorem 8).
-An $((n,K,d))_{\mathfrak{q}}$ stabilizer code exists if and only if there exists an additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$ if $K>1$, and $\operatorname{swt}(C^{\perp_{\chi}}\backslash\{0\})=d$ if $K=1$.
-A linear code over the ring $R$ of length $n$, rank $k$, and minimum Hamming distance $d$ is denoted by the parameters $[n,k,d]_{R}$.
-Now, we are ready to present an annihilator counterpart for the Euclidean and Hermitian CSS construction, as presented in [17, 29, 33].
+The Euclidean CSS construction is applied to classical linear codes that preserve Euclidean duals to get quantum codes, see for example [34].
+In this section, we aim to introduce the annihilator CSS construction for polycyclic codes based on the annihilator dual.
+Let $R$ be a finite commutative Frobenius ring with $\mathfrak{q}$ elements.
+Let $f(x)=x^{n}-\sum_{r=0}^{n-1}f_{r}x^{r}\in R[x]$ and $C$ be an $f(x)$-polycyclic code.
+According to Section 6, the annihilator product $\langle g(x),h(x)\rangle_{f}$ is defined as $r(0)$, where $r(x)$ is the reminder of $gh(x)$ divided by $f$.
+The annihilator dual with respect to this product is denoted by $C^{\circ}.$ The Gram matrix $A$ associated with the polynomial $f$ is a $n\times n$ matrix with entries $a_{i,j}=\langle x^{i-1},x^{j-1}\rangle_{f}$.
+For example, if $f(x)=x^{5}-f_{4}x^{4}-f_{3}x^{3}-f_{2}x^{2}-f_{1}x-f_{0}$, then
+$A=\begin{pmatrix}1&0&0&0&0\\[4.0pt] 0&0&0&0&f_{0}\\[4.0pt] 0&0&0&f_{0}&f_{0}f_{4}\\[4.0pt] 0&0&f_{0}&f_{0}f_{4}&f_{0}(f_{4}^{2}+f_{3})\\[4.0pt] 0&f_{0}&f_{0}f_{4}&f_{0}(f_{4}^{2}+f_{3})&f_{0}\left(f_{4}^{3}+2f_{4}f_{3}+f_{2}\right)\end{pmatrix}.$ (21)
+The Gram matrix $A$ is symmetric, and if $f_{0}$ is invertible, then $A$ is also invertible. The following lemma provides a recursive formula for calculating the matrix $A$.
+Lemma 3.
+Let $f(x)=x^{n}-\sum_{r=0}^{n-1}f_{r}x^{r}\in R[x]$ and $A=(a_{i,j})$ be an $n\times n$ Gram matrix associated with $f$. Then $a_{i,j}\;=\;c_{\,i+j-2}$, where
+$c_{k}=\begin{cases}1,&k=0,\\ 0,&0<k<n,\\ \sum_{r=0}^{n-1}f_{r}\,c_{k-n+r},&k\geq n.\end{cases}$
+Proof.
+To compute each $a_{i,j}$, it suffices to find the constant term of the associated polynomial
+$x^{\,i+j-2}\bigr{\rvert}_{x^{n}=f_{n-1}x^{n-1}+\cdots+f_{1}x+f_{0}}.$
+Let us denote the constant term of $x^{k}|_{x^{n}}$ by $c_{k}$.
+Therefore, $A$ is the following Hankel matrix [^1] A matrix is called Hankel if each ascending skew-diagonal from left to right is constant.
+$A=\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-2}&c_{n-1}\\ c_{1}&c_{2}&c_{3}&\ldots&c_{n-1}&c_{n}\\ c_{2}&c_{3}&c_{4}&\ldots&c_{n}&c_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ c_{n-2}&c_{n-1}&c_{n}&\ldots&c_{2n-4}&c_{2n-3}\\ c_{n-1}&c_{n}&c_{n+1}&\ldots&c_{2n-3}&c_{2n-2}\end{pmatrix}.$
+It is clear that $c_{k}=1$ if $k=0$, and $c_{k}=0$ if $0<k<n$. For $k\geqslant n$, we have $x^{k}=x^{k-n}x^{n}=x^{k-n}\sum_{r=0}^{n-1}f_{r}\,x^{r}=\sum_{r=0}^{n-1}f_{r}\,x^{(k-n)+r}$. Therefore,
+$c_{k}=\left(\text{constant term of }x^{k}\bigr{\rvert}_{x^{n}}\right)=\sum_{r=0}^{n-1}f_{r}\left(\left.\text{constant term of }x^{(k-n)+r}\right|_{x^{n}}\right)=\sum_{r=0}^{n-1}f_{r}\,c_{\,k-n+r}.$
+Lemma 4.
+If $R$ is a commutative ring and $f_{0}$ is invertible, then $C^{\circ}=(CA)^{\perp}=C^{\perp}A^{-1}$, where $CA=\{cA\,|\,c\in C\}$.
+Proof.
+[…]
+Remark 6.
+Suppose that $f(x)=x^{n}-f_{0}\in R[x]$ with $f_{0}$ invertible. Then $A$ and its inverse are the monomial matrices
+$A=\begin{pmatrix}1&0&\ldots&0&0\\ 0&0&\ldots&0&f_{0}\\ 0&0&\ldots&f_{0}&0\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&f_{0}&\ldots&0&0\end{pmatrix},\qquad A^{-1}=\begin{pmatrix}1&0&\ldots&0&0\\ 0&0&\ldots&0&f_{0}^{-1}\\ 0&0&\ldots&f_{0}^{-1}&0\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&f_{0}^{-1}&\ldots&0&0\end{pmatrix}.$
+In this case, since $C^{\circ}=C^{\perp}A^{-1}$, the dual codes $C^{\circ}$ and $C^{\perp}$ are monomially equivalent, which ensures they have the same minimum Hamming distance.
+However, in general, $A$ may not preserve the minimum Hamming distance because it may not preserve the Hamming weight.
+Below is a well-known theorem for constructing quantum codes from classical linear codes that preserve the Euclidean dual.
+Note that $[[n,k,d]]_{\mathfrak{q}}$ will denote a $R$-ary quantum code of length $n$, with $\mathfrak{q}^{k}$ codewords, and the minimum distance $d$.
+Theorem 13 (Euclidean CSS construction [34]).
+Let $C_{1}$ and $C_{2}$ be two linear codes over the finite commutative Frobenius ring $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively.
+If $C_{2}^{\perp}\subseteq C_{1}$, then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\perp})\cup(C_{2}\setminus C_{1}^{\perp})\}$ is the minimum Hamming distance.
+Now, we are ready to introduce the CSS construction for polycyclic codes that preserve annihilator duals.
 Theorem 14 (Annihilator CSS construction).
-Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$.
-Let $f_{0}\in R$ be an invertible element, $f(x)=g_{1}(x)h_{1}(x)=g_{2}(x)h_{2}(x)$, $C_{1}=\langle g_{1}(x)\rangle$ and $C_{2}=\langle g_{2}(x)\rangle$ be two $f(x)$-polycyclic codes over the ring $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively, such that $C^{\circ}_{2}\subseteq C_{1}$.
-Then, there exists an $((n,k_{1}+k_{2}-n,d))_{\mathfrak{q}}$ stabilizer code with the minimum distance $d=\min\{\operatorname{wt}(c)\,|\,c\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})\}$.
+Suppose that $R$ is a Frobenius ring and $f_{0}\in R$ is invertible.
+Let $f(x)=g_{1}(x)h_{1}(x)=g_{2}(x)h_{2}(x)$, and let $C_{1}=\langle g_{1}(x)\rangle$ and $C_{2}=\langle g_{2}(x)\rangle$ be two $f(x)$-polycyclic codes over $R$ with parameters $[n,k_{1},d_{1}]_{R}$ and $[n,k_{2},d_{2}]_{R}$, respectively.
+If $C^{\circ}_{2}\subseteq C_{1}$ then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\circ})\cup(C_{2}A\setminus C_{1}^{\circ}A)\}$ is the minimum Hamming distance.
 Proof.
-We define $C=C^{\circ}_{1}\times C^{\circ}_{2}\subseteq{R}^{2n}$, where $C_{1}^{\circ}$ and $C_{2}^{\circ}$ are annihilator duals of $C_{1}$ and $C_{2}$, respectively.
-If $(a|b)$ and $(a^{\prime}|b^{\prime})$ are in $C$, then $b\cdot a^{\prime}-b^{\prime}\cdot a=0$.
-Thus $C\subseteq C^{\perp_{s}}$, that is $C\perp_{s}C$.
-Using Lemma 4, we have $C\subseteq C^{\perp_{\chi}}$.
-Furthermore, if $(a|b)\in C_{2}\times C_{1}$, we have $b\cdot a^{\prime}-b^{\prime}\cdot a=0$ for any arbitrary element $(a^{\prime}|b^{\prime})$ in $C$, and hence $(a|b)\in C^{\perp_{s}}$.
-So, $(C_{2}\times C_{1})\perp_{s}C$.
-Applying Lemma 4, $(C_{2}\times C_{1})\perp_{\chi}C$, which implies $(C_{2}\times C_{1})\subseteq C^{\chi}$.
-On the other side, by Lemma 3 and Part (3) in Theorem 10, we have
-$|C^{\perp_{\chi}}|=\frac{|R|^{2n}}{|C|}=\frac{|R|^{n}}{|C^{\circ}_{1}|}\frac{|R|^{n}}{|C^{\circ}_{2}|}=|C_{1}||C_{2}|=|C_{2}\times C_{1}|.$
-Therefore, $C_{2}\times C_{1}=C^{\perp_{\chi}}$. Moreover,
-$|C|=|C^{\circ}_{1}||C^{\circ}_{2}|=\frac{|R|^{n}}{|C_{1}|}\frac{|R|^{n}}{|C_{2}|}=\frac{|R|^{2n}}{|R|^{k_{1}+k_{2}}}=\frac{|R|^{n}}{|R|^{k_{1}+k_{2}-n}}.$
-If we take $K=|{R}|^{k_{1}+k_{2}-n}$, then $|C|=|R|^{n}/K$.
-We claim that $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)=d$.
-To prove, assume that $(a|b)\in(C^{\perp_{\chi}}\backslash C)=((C_{2}\times C_{1})\backslash(C^{\circ}_{1}\times C^{\circ}_{2}))$ is an arbitrary element.
-So, $(a|b)\in C_{2}\times C_{1}$ and $(a|b)\notin C^{\circ}_{1}\times C^{\circ}_{2}$.
-That is $a\notin C^{\circ}_{1}$ or $b\notin C^{\circ}_{2}$.
-Therefore, $a,b\in(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$.
-According to the definition of $d$, $\operatorname{wt}(a)\geqslant d$ and $\operatorname{wt}(b)\geqslant d$.
-Applying Equation (21), we have
-$\operatorname{swt}(a|b)\geqslant\frac{1}{2}\operatorname{wt}(a|b)=\frac{1}{2}(\operatorname{wt}(a)+\operatorname{wt}(b))\geqslant d.$
-Thus, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\geqslant d$.
-On the other side, we know there exists an element $c$ in $(C_{1}\backslash C_{2}^{\circ})\cup(C_{2}\backslash C_{1}^{\circ})$ such that $\operatorname{wt}(c)=d$.
-In the first case, let $c\in(C_{1}\backslash C_{2}^{\circ})$.
-Then we have $(0|c)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(0|c)=\operatorname{wt}(c)=d$.
-In the second case, let $c\in(C_{2}\backslash C_{1}^{\circ})$.
-Then we have $(c|0)\in(C^{\perp_{\chi}}\backslash C)$, and $\operatorname{swt}(c|0)=\operatorname{wt}(c)=d$.
-Hence in both cases, there exists an element in $(C^{\perp_{\chi}}\backslash C)$ with the symplectic weight $d$.
-Therefore, $\operatorname{swt}(C^{\perp_{\chi}}\backslash C)\leqslant d$.
-To sum up, we have found the additive code $C\subseteq R^{2n}$ of size $|C|=|R|^{n}/K$ such that $C\subseteq C^{\perp_{\chi}}$ and $\operatorname{wt}(C^{\perp_{\chi}}\backslash C)=d$ As a result, Theorem 13 implies that there exists an $((n,K,d))_{\mathfrak{q}}$ stabilizer code.
+Since the $n\times n$ Gram matrix $A$ is invertible over $R$, then $xA=0$ leads to $x=0$ for any $x\in R^{n}$ (similar to Corollary 2.4 in [27]).
+So the transformation $x\mapsto xA$ is a bijective over $R^{n}$, which implies that the dimension of the code $C_{2}A$ equals to $k_{2}$.
+So, if we replace $C_{2}$ by $C_{2}A$ in the Theorem above, then there exists a quantum stabilizer code with parameters $[[n,k_{1}+k_{2}-n,d]]_{\mathfrak{q}}$, where $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus(C_{2}A)^{\perp})\cup(C_{2}A\setminus C_{1}^{\perp})\}$.
+Applying Lemma 4, we obtain $d=\min\{\operatorname{wt}(c)\mid c\in(C_{1}\setminus C_{2}^{\circ})\cup(C_{2}A\setminus C_{1}^{\circ}A)\}$.
 ∎
 Corollary 4.
-Suppose that $R$ is a chain ring and $f(x)$ defined as in Equation (3) is a polynomial in $R[x]$.
-Let $f_{0}\in R$ be an invertible element and $f(x)=g(x)h(x)$.
-If $C$ is an $f(x)$-polycyclic code over the chain ring $R$ with parameters $[n,k,d]_{R}$, generated by $g(x)$, and $C^{\circ}\subseteq C$, then there exists an $((n,2k-n,\geqslant d))_{\mathfrak{q}}$ stabilizer code.
-Now, we are prepared to give the result that allows to construct quantum codes from $f(x)$-polycyclic codes over the serial ring $\mathscr{R}$.
-Theorem 15.
-Suppose that $f(x)$ defined as in Equation (3) is a polynomial in $\mathscr{R}[x]$, $f_{0}\in\mathscr{R}$ is an invertible element, and $f(x)=g(x)h(x)$.
-Let $\mathscr{C}$ be an annihilator dual-preserving $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$, decomposed as $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}$.
-If each $\mathscr{C}_{i}$ has parameters $[n_{i},k_{i},d_{i}]$, then there exists a stabilizer code with parameters $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$.
-Proof.
-By Corollary 3, $\mathscr{C}_{i}$ is dual-preserving $f_{i}(x)$-polycyclic code over the chain ring $\mathscr{R}_{i}$ for all $i$.
-Suppose that $\big{(}g_{1}(x),g_{2}(x),\ldots,g_{k}(x)\big{)}$ and $\big{(}h_{1}(x),h_{2}(x)\ldots,h_{k}(x)\big{)}$ are the images of $g(x)$ and $h(x)$ under the mapping $\eta$ defined in Eq.
-(7), respectively.
-Since $g(x)=\oplus_{i=1}^{k}g_{i}(x)e_{i}$ and $h(x)=\oplus_{i=1}^{k}h_{i}(x)e_{i}$, we have $f_{i}(x)=g_{i}(x)h_{i}(x)$ for all $i$.
-Moreover, since $f_{0}$ is invertible, the constant term of the polynomial $f_{i}(x)$ is invertible for all $i$.
-By Corollary 4, there exists an $((n_{i},2k_{i}-n_{i},\geqslant d_{i}))$ stabilizer code for all $i$.
-It is well-know that the direct sum of two quantum-error-correcting codes $((n,k,d))_{\mathfrak{q}}$ and $((n^{\prime},k^{\prime},d^{\prime}))_{\mathfrak{q}}$ is an $((n+n^{\prime},k+k^{\prime},d^{\prime\prime}))_{\mathfrak{q}}$ code, where $d^{\prime\prime}=\min\{d,d^{\prime}\}$, see [8] Section 4.
-Therefore, there exists a $((\sum_{i=1}^{k}n_{i},\sum_{i=1}^{k}(2k_{i}-n_{i}),\geqslant d))_{\mathfrak{q}}$ stabilizer code, where $d\geqslant\min\{d_{1},d_{2},\ldots,d_{k}\}$.
-∎
-8 Conclusion
+Suppose that $R$ is a finite Frobenius ring and $f_{0}\in R$ is invertible.
+Let $f(x)=g(x)h(x)$, and let $C=\langle g(x)\rangle$ be an $f(x)$-polycyclic code over $R$ with parameters $[n,k,d_{1}]_{R}$.
+If $C^{\circ}\subseteq C$, then there exists an $[[n,2k-n,d]]_{\mathfrak{q}}$ stabilizer code, where $d=\min\{\operatorname{wt}(c)\mid c\in(C\setminus C^{\circ})\cup(CA\setminus C^{\circ}A)\}$.
+Example 7.1.
+Consider the serial ring $\mathscr{R}=\mathbb{F}_{3}[x_{1},x_{2}]/\langle x_{1}^{2}-x_{1},x_{2}^{2}-x_{2}\rangle$, which is isomorphic to $\mathbb{F}_{3}+u\mathbb{F}_{3}+v\mathbb{F}_{3}+uv\mathbb{F}_{3}$ with $u^{2}=u$, $v^{2}=v$, and $uv=vu$.
+According to Proposition 2, $\mathscr{R}$ is a Frobenius ring.
+Let $f(x)=x^{5}+2x^{3}+x^{2}+2=(x+1)^{4}(x+2)\in\mathscr{R}[x]$.
+The $f(x)$-polycyclic code $\mathscr{C}$ generated by $g(x)=x+1$ is a free $\mathscr{R}$-module of dimension $k=\deg(f)-\deg(g)$ with the generating set $\{g(x),xg(x),\ldots,x^{k-1}g(x)\}$.
+Consider the Gray map $\varphi:\mathscr{R}\rightarrow\mathbb{F}_{3}^{4}$ defined as $\varphi(a+ub+vc+uvd)=(a,a+b,a+c,a+b+c+d)$ in [5].
+Using equation (21), one can compute the Gram matrix $A$.
+Aplying Theorem 12, $\mathscr{C}$ is annihilator dual-containing.
+Using Magma [11], we compute that $d=2$.
+Therefore, we obtain a $[[5,3,2]]_{3^{4}}$ quantum stabilizer code.
+8 Conclusion and future work
 In this paper, we have investigated the structure of $f(x)$-polycyclic codes over a specific category of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial.
@@
 We have defined the new concepts QsDP codes and $f(x)$-polycyclic-QsDP codes, and have proved that $f(x)$-polycyclic codes over $\mathscr{R}$ are isomorphic to $f(x)$-polycyclic-QsDP codes.
-Additionally, we have established a polynomial structure for the Eclidean dual of $f(x)$-polycyclic codes over $\mathscr{R}$.
-The annihilator dual is an alternative duality concept ensuring that the dual of $f(x)$-polycyclic codes has an $f(x)$- polycyclic structure, a feature that fails in Euclidean dual.
+Additionally, we have established a polynomial structure for the Euclidean dual of $f(x)$-polycyclic codes over $\mathscr{R}$.
+The annihilator dual is an alternative duality concept ensuring that the dual of $f(x)$-polycyclic codes have an $f(x)$- polycyclic structure, a feature that fails in the Euclidean dual.
 We have studied some properties of the annihilator dual of $f(x)$-polycyclic code over the ring $\mathscr{R}$ and have provided necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator LCD, and annihilator dual-containing polycyclic codes over $\mathscr{R}$.
-Finally, we have provided an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.
+Finally, we have provided an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes over Frobenius rings.
+The use of a general polynomial other than $x^{n}-1$ or $x^{n}-\lambda$ (cyclic and constacyclic case) provides the chance to explore a broader range of factors, and the annihilator duality provides that the dual code is always in the same ambient space and it is easy to compute, hence the annihilator CSS construction provides a broader class of codes where to find quantum codes.
+As a future work, it will be worth studying the annihilator duality and checking if it yields new quantum codes.
+Also in [7], the authors have recently studied quasi-polycyclic codes and their application for constructing quantum error-correcting codes based on the CSS construction for Euclidean and symplectic duality.
+Extending these results to the annihilator dual setting may yield new classes of quantum codes.